### v1

# Counting rational points on Hirzebruch–Kleinschmidt varieties over global function fields

Abstract.
Inspired by Bourqui’s work on anticanonical height zeta functions on Hirzebruch surfaces, we study height zeta functions of split toric varieties with Picard rank 2 over global function fields, with respect to height functions associated with big metrized line bundles. We show that these varieties can be naturally decomposed into a finite disjoint union of subvarieties, where precise analytic properties of the corresponding height zeta functions can be given, allowing for a finer inspection of the asymptotic number of rational points of bounded height on each subvariety.
2020 Mathematics Subject Classification:
14G05, 11G50, 11M41 (primary), 14M25, 11G35 (secondary)
1. Introduction
In the papers [FMT89] and [BM90], a program was initiated to understand the asymptotic behavior of the number of rational points of bounded height on Fano algebraic varieties over global fields. In the case of number fields, the current expectation is as follows (see e.g. [Pey03, Formule empirique 5.1] or [ADHL15, Conjecture 6.3.1.5] for precise definitions).
Conjecture 1.1 (Manin–Peyre).
Let $X$ be an almost Fano variety [^1] Following [Pey03, Définition 3.1], an almost Fano variety is a smooth, projective, geometrically integral variety $X$ defined over a field $K$, with $H^{1}(X,\mathscr{O}_{X})=H^{2}(X,\mathscr{O}_{X})=0$, torsion-free geometric Picard group $\operatorname{Pic}(X_{\overline{K}})$ and $-K_{X}$ big. over a number field $K$, with dense set of rational points $X(K)$, finitely generated effective cone  $\Lambda_{\textup{eff}}(X_{\overline{K}})$ and trivial Brauer group $\operatorname{Br}(X_{\overline{K}})$. Let $H=H_{-K_{X}}$ be the anticanonical height function, and assume that there is an open subset $U$ of $X$ that is the complement of the weakly accumulating subvarieties on $X$ with respect to $H$. Then, there is a constant $C>0$ such that
$N(U,H,B):=\#\{P\in U(K):H(P)\leq B\}\sim CB(\log B)^{\operatorname{rk}\operatorname{Pic}(X)-1}\quad\text{as }B\rightarrow\infty.$ (1)
Moreover, the leading constant is of the form
$C=\alpha(X)\beta(X)\tau_{H}(X),$
where
$\alpha(X)$ $:=\frac{1}{\left(\operatorname{rk}\operatorname{Pic}(X)-1\right)!}\int_{\Lambda_{\textup{eff}}(X)^{\vee}}e^{-\left\langle-K_{X},\mathbf{y}\right\rangle}\textup{d}\mathbf{y},$
$\beta(X)$ :=#$:=\#H^{1}(\textup{Gal}(\overline{K}/K),\operatorname{Pic}(X_{\overline{K}})),$
and $\tau_{H}(X)$ is the Tamagawa number of $X$ with respect to $H$ defined by Peyre in [Pey03, Section 4].
There is also a conjecture, originating in the work of Batyrev and Manin [BM90], concerning the asymptotic growth of the number
$N(U,H_{L},B)=\#\{P\in U(K):H_{L}(P)\leq B\},$
when considering height functions $H_{L}$ associated with big metrized line bundles $L$ on a variety $X$ as above, and for appropriate open subsets $U\subseteq X$. More precisely, if we write $\tau\prec\sigma$ whenever $\tau$ is a face of a cone $\sigma$, then one defines the following classical numerical invariants for a line bundle class $L$ on $X$:
$\begin{split}a(L)&:=\inf\{a\in\mathbb{R}:aL+K_{X}\in\Lambda_{\textup{eff}}(X)\},\\ b(L)&:=\max\{\textup{codim}(\tau):a(L)L+K_{X}\in\tau\prec\Lambda_{\textup{eff}}(X)\},\end{split}$
which measure the position of $L$ inside the effective cone $\Lambda_{\textup{eff}}(X)$. With this notation, the more general version of the above conjecture states that there exists a constant $C>0$ such that
$N(U,H_{L},B)\sim CB^{a(L)}(\log B)^{b(L)-1}\quad\text{as }B\to\infty.$ (2)
Note that $a(-K_{X})=1$ and $b(-K_{X})=\operatorname{rk}\operatorname{Pic}(X)$, hence (2) is indeed a generalization of (1).
The above conjectures have been proven by various authors, either in specific examples or in certain families of varieties (see for instance [Pey03], [Tsc03, Section 3], [Bro07] and [Tsc09, Section 4] for accounts of such results).
Typically, asymptotic formulas of the form (1) and (2) are deduced, via a Tauberian theorem, from the analytic properties of the associated height zeta functions. More precisely, considering $U\subseteq X$ and a height function $H_{L}$ associated to a big metrized line bundle $L$, the corresponding height zeta function is defined as
$\zeta_{U,L}(s):=\sum_{P\in U(K)}H_{L}(P)^{-s}.$
If $\zeta_{U,L}(s)$ converges absolutely in the half-plane $\Re(s)>a>0$, has analytic continuation to $\Re(s)>a-\varepsilon,s\neq a$ for some $\varepsilon>0$ and has a pole of order $b\geq 1$ at $s=a$, then one obtains
$N(U,H_{L},B)\sim CB^{a}(\log B)^{b-1}\quad\text{as }B\to\infty,$ (3)
with
$C:=\frac{1}{(b-1)!a}\lim_{s\to a}(s-a)^{b}\zeta_{U,L}(s)$
(see e.g. [Del54, Théorème III]).
In the case where $K$ is a global function field of positive characteristic, the relevant height functions typically have values contained in $q^{\mathbb{Z}}$, where $q$ the cardinality of the constant subfield $\mathbb{F}_{q}\subset K$. This implies that the associated height zeta functions are invariant under $s\mapsto s+\frac{2\pi i}{\log(q)}$. In particular, having a pole at a point $s=a>0$ implies the existence of infinitely many poles on the line $\Re(s)=a$, which makes it impossible to apply a Tauberian theorem. Moreover, in this setting, it is hopeless to expect an asymptotic formula of the form (3) since $N(U,H_{L},q^{n})=N(U,H_{L},q^{n+\frac{1}{2}})$ would imply $\sqrt{q}=1$ (as explained in [Bou11, Section 1.1]).
For the reasons explained above, for varieties defined over global fields of positive characteristic, the analogue of Conjecture 1.1 is presented in terms of the analytical properties of the respective height zeta function. This is the approach taken by Peyre in [Pey12] in the case of flag varieties. As a variant of Conjecture 1.1, it is then expected that the anticanonical height zeta function $\zeta_{U,-K_{X}}(s)$, for appropriate open subsets $U\subseteq X$, converges absolutely for $\Re(s)>1$, it admits meromorphic continuation to $\Re(s)>1-\varepsilon$ for some $\varepsilon>0$, and has a pole at $s=1$ of order $\operatorname{rk}\operatorname{Pic}(X)$ satisfying
$\lim_{s\to 1}(s-1)^{\operatorname{rk}\operatorname{Pic}(X)}\zeta_{U,-K_{X}}(s)=\alpha^{\ast}(X)\beta(X)\tau_{H}(X),$
where
$\alpha^{\ast}(X)$ $:=\int_{\Lambda_{\textup{eff}}(X)^{\vee}}e^{-\left\langle-K_{X},\mathbf{y}\right\rangle}\textup{d}\mathbf{y},$
$\beta(X)$ :=#$:=\#H^{1}(\textup{Gal}(K^{\mathrm{sep}}/K),\operatorname{Pic}(X_{K^{\mathrm{sep}}})),$
and $\tau_{H}(X)$ is the Tamagawa number of $X$ with respect to the anticanonical height function $H=H_{-K_{X}}$ defined by Peyre in [Pey12, Section 2] in the case of global fields of positive characteristic. Similarly, as a variant of (2), one expects $\zeta_{U,L}(s)$ to converge absolutely for $\Re(s)>a(L)$ and have meromorphic continuation to $\Re(s)>a(L)-\varepsilon$ for some $\varepsilon>0$ with a pole at $s=a(L)$ of order $b(L)$.
In this paper we focus on smooth projective split toric varieties with Picard rank 2 over a global field $K$ of positive characteristic, and consider height zeta functions associated to big metrized line bundles. Our results show that these varieties can be naturally decomposed into a finite disjoint union of subvarieties for which precise analytic properties of the corresponding height zeta functions can be given. Hence, in this setting, we go beyond the scope of the classical expectations mentioned above. We achieve these results by using concrete defining equations for such varieties and by performing explicit computations on the height zeta functions.
The present work is motivated by the results of Bourqui in [Bou02], [Bou03] and [Bou11], where the above expectation for the anticanonical height zeta function is verified for toric varieties over global fields of positive characteristic, with $U$ the dense torus orbit. Bourqui’s work was inspired by Batyrev and Tschinkel’s proof of Conjecture 1.1 for toric varieties over number fields [BT98].
In order to present our results more precisely, let us recall a geometric result due to Kleinschmidt [Kle88] stating that all smooth projective toric varieties of Picard rank 2 are (up to isomorphism) of the form
$X=\mathbb{P}(\mathscr{O}_{\mathbb{P}^{t-1}}\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(a_{1})\oplus\cdots\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(a_{r})),$
where $r\geq 1$, $t\geq 2$, and $0\leq a_{1}\leq\cdots\leq a_{r}$ are integers. Moreover, if we define $I_{k}:=\{1,\ldots,k\}\subset\mathbb{Z}$, then $X$ is isomorphic to the subvariety $X_{d}(a_{1},\ldots,a_{r})\subset\mathbb{P}^{rt}\times\mathbb{P}^{t-1}$ given in homogeneous coordinates $([x_{0}:(x_{ij})_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t}])$ by the equations
$x_{mj}y^{a_{j}}_{n}=x_{nj}y^{a_{j}}_{m},\,\text{ for all }j\in I_{r}\text{ andall }m,n\in I_{t}\text{ with }m\neq n.$
We refer to these varieties as Hirzebruch–Kleinschmidt varieties. As usual, in order to avoid accumulation of rational points of bounded height, we find it necessary to restrict our attention to rational points in a dense open subset. To this purpose, we define the good open subset $U_{d}(a_{1},\ldots,a_{r})\subset X_{d}(a_{1},\ldots,a_{r})$ as the complement of the closed subvariety define by the equation $x_{tr}=0$.
Our main results describe the analytic properties of the height zeta function $\zeta_{U,L}(s)$ for the good open subset $U:=U_{d}(a_{1},\ldots,a_{r})\subset X_{d}(a_{1},\ldots,a_{r})$, for every big line bundle class $L\in\operatorname{Pic}(X)$ that we equip with a “standard metrization”. These general results are given in Section 5.1. For simplicity, we present here the statement for $L=-K_{X}$.
We let $\mathscr{C}$ be a projective, smooth, geometrically irreducible curve of genus $g$ defined over the finite field $\mathbb{F}_{q}$ of $q$ elements, and let $K=\mathbb{F}_{q}(\mathscr{C})$ denote the field of rational functions of $\mathscr{C}$.
Theorem 1.2.
Let $X=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety of dimension $d=r+t-1$ over a global function field $K=\mathbb{F}_{q}(\mathscr{C})$, let $H=H_{-K_{X}}$ denote the anticanonical height function on $X(K)$, and let $U=U_{d}(a_{1},\ldots,a_{r})$ be the good open subset of $X$. Then, the anticanonical height zeta function $\zeta_{U,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$, and it is a rational function in $q^{-s}$. In particular, it has meromorphic continuation to $\mathbb{C}$. Moreover, it has a pole of order 2 at $s=1$ with
$\lim_{s\to 1}(s-1)^{2}\zeta_{U,-K_{X}}(s)=\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)((r+1)a_{r}-|\mathbf{a}|+t)(r+1)(q-1)^{2}\log(q)^{2}},$ (4)
where $|\mathbf{a}|:=\sum_{i=1}^{r}a_{i}$, $h_{K}$ is the class number of $K$, $g$ is the genus of $\mathscr{C}$ and $\zeta_{K}$ is the zeta function of $K$.
Remark 1.3.
In Section 3.2 we show that
$\alpha^{\ast}(X)=\frac{1}{(r+1)((r+1)a_{r}-|\mathbf{a}|+t)}.$
Also, since $X$ is a split toric variety, it is well known that $\beta(X)=1$ (see e.g. [BT98, Remark 1.7 and Corollary 1.18]). Hence, Theorem 1.2 and [Bou03, Théorème 1.1] imply that the Tamagawa number of $X$ with respect to the anticanonical height function $H=H_{-K_{X}}$ is
$\tau_{H}(X)=\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(q-1)^{2}\log(q)^{2}}.$
As mentioned before, we go beyond the classical expectations and study further the height zeta function of the complement of the good open subset inside $X_{d}(a_{1},\ldots,a_{r})$. This is achieved by a natural decomposition of $X=X_{d}(a_{1},\ldots,a_{r})$ of the form
$X\simeq\left\{\begin{array}[]{ll}X_{d-1}(a_{1},\ldots,a_{r-1})\sqcup\mathbb{A}^{r}\sqcup\bigg{(}\bigsqcup_{2\leq t^{\prime}\leq t}U_{t^{\prime}+r-1}(a_{1},\ldots,a_{r})\bigg{)}&\text{if }r>1,\\ \mathbb{P}^{t-1}\sqcup\mathbb{A}^{1}\sqcup\bigg{(}\bigsqcup_{2\leq t^{\prime}\leq t}U_{t^{\prime}}(a_{1})\bigg{)}&\text{if }r=1\end{array}\right.$ (5)
(see Section 5.4 for details), which allows us to work recursively on the dimension of $X$, and decompose $\zeta_{X,-K_{X}}(s)$ as a finite sum of height zeta functions of projective (and affine) spaces and height zeta functions of good open subsets of Hirzebruch–Kleinschmidt varieties of dimension $d=t+r-1,\ldots,2$. On the one hand, height zeta functions of projective (and affine) spaces are well understood (see Section 4). On the other hand, for each good open subset $U^{\prime}=U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})\subseteq U_{d}(a_{1},\ldots,a_{r})$ with $2\leq t^{\prime}\leq t,1\leq r^{\prime}\leq r$ we get a contribution of a height zeta function of the form $\zeta_{U^{\prime},L}(s)$ with $L\in\operatorname{Pic}(X_{d^{\prime}}(a_{1},\ldots,a_{r^{\prime}}))$ generally distinct from the corresponding anticanonical class. Nevertheless, provided $L$ is big, we can still describe the analytic properties of $\zeta_{U^{\prime},L}(s)$ in detail, see Theorems 5.2 and 5.4 in Section 5.1. Note that the bigness condition on $L$ is necessary, because when $L$ is not big, $\zeta_{U^{\prime},L}(s)$ has no finite abscissa of absolute convergence. The following example illustrates these ideas.
Example 1.4.
Given an integer $a>0$, let us consider the Hirzebruch–Kleinschmidt threefold $X:=X_{3}(a)$ with projection map $\pi:X_{3}(a)\to\mathbb{P}^{2}$. The decomposition (5) in this case becomes
$X\simeq\mathbb{P}^{2}\sqcup\mathbb{A}^{1}\sqcup U_{3}(a)\sqcup U_{2}(a).$ (6)
Put $U:=U_{3}(a),U^{\prime}:=U_{2}(a)\subset X^{\prime}:=X_{2}(a)$ and let $\pi^{\prime}:X^{\prime}\to\mathbb{P}^{1}$ denote the corresponding projection map. The anticanonical class on $X$ is $-K_{X}=\mathscr{O}_{X}(2)\otimes\pi^{\ast}(\mathscr{O}_{\mathbb{P}^{2}}(3-a))$. Using Theorem 1.2 we get that $\zeta_{U,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$ and has meromorphic continuation to $\mathbb{C}$ with a pole of order 2 at $s=1$ with
$\lim_{s\to 1}(s-1)^{2}\zeta_{U,-K_{X}}(s)=\frac{q^{5(1-g)}h_{K}^{2}}{\zeta_{K}(3)\zeta_{K}(2)(a+3)2(q-1)^{2}\log(q)^{2}}.$
By Lemma 5.15 in Section 5.4 the component $U_{2}(a)$ in (6) contributes with the height zeta function $\zeta_{U^{\prime},L}(s)$ with $L:=\mathscr{O}_{X^{\prime}}(2)\otimes(\pi^{\prime})^{\ast}(\mathscr{O}_{\mathbb{P}^{1}}(3-a))\in\operatorname{Pic}(X^{\prime})$, which is big but different from the anticanonical class $-K_{X^{\prime}}$. Nevertheless, by Theorem 5.2 we know that $\zeta_{U^{\prime},L}$ converges absolutely for $\Re(s)>1$ and has meromorphic continuation to $\mathbb{C}$ with a simple pole at $s=1$.
Now, by the same Lemma 5.15 the component $\mathbb{P}^{2}$ in (6) contributes with the height zeta function $\zeta_{\mathbb{P}^{2}}((3-a)s)$ where $\zeta_{\mathbb{P}^{2}}(s)$ denotes the “standard height zeta function” of $\mathbb{P}^{2}$ defined in Section 4. We see that $\zeta_{\mathbb{P}^{2}}((3-a)s)$ has no finite abscissa of absolute convergence when $a\geq 3$. Assuming $a<3$, we can apply Theorem 4.1 in Section 4 to deduce that $\zeta_{\mathbb{P}^{2}}((3-a)s)$ converges absolutely for $\Re(s)>\frac{3}{3-a}$ and has meromorphic continuation to $\mathbb{C}$ with a simple pole at $s=\frac{3}{3-a}>1$.
Finally, by Lemma 5.15 the component $\mathbb{A}^{1}$ in (6) contributes with the height zeta function $\zeta_{\mathbb{A}^{1}}(2s)=\zeta_{\mathbb{P}^{1}}(2s)-1$, where $\zeta_{\mathbb{P}^{1}}(s)$ denotes the standard height zeta function of $\mathbb{P}^{1}$. In particular, by Theorem 4.1 we know that $\zeta_{\mathbb{A}^{1}}(2s)$ converges absolutely for $\Re(s)>1$ and has meromorphic continuation to $s\in\mathbb{C}$ with a simple pole at $s=1$.
We can then distinguish the following two cases:
  1. (1)
    If $a=1,2$, then the number of rational points of bounded anticanonical height in $X_{3}(a)$ is “dominated” by the number of those points in the component $\mathbb{P}^{2}$ in (6). More precisely, we have that
    $\zeta_{X,-K_{X}}(s)=\zeta_{\mathbb{P}^{2}}((3-a)s)+\zeta_{\mathbb{A}^{1}}(2s)+\zeta_{U,-K_{X}}(s)+\zeta_{U^{\prime},L}(s)$
    converges absolutely for $\Re(s)>\sigma_{a}:=\frac{3}{3-a}>1$ and has meromorphic continuation to $\mathbb{C}$ with a simple pole at $s=\sigma_{a}$, while
    $\zeta_{X\setminus\mathbb{P}^{2},-K_{X}}(s)=\zeta_{\mathbb{A}^{1}}(2s)+\zeta_{U,-K_{X}}(s)+\zeta_{U^{\prime},L}(s)$
    converges absolutely for $\Re(s)>1$. Moreover, by Theorem 4.1 in Section 4, we can compute
    $\lim_{s\to\sigma_{a}}\left(s-\sigma_{a}\right)\zeta_{X,-K_{X}}(s)=\lim_{s\to\sigma_{a}}\left(s-\sigma_{a}\right)\zeta_{\mathbb{P}^{2}}((3-a)s)=\frac{h_{K}q^{3(1-g)}}{(3-a)\zeta_{K}(3)(q-1)\log(q)}.$
  2. (2)
    If $a\geq 3$, then the component $\mathbb{P}^{2}$ in (6) has “too many” rational points of bounded height. In fact, $N(\mathbb{P}^{2},-K_{X},B)=\infty$ for all $B\geq 1$. In this case we also have
    $\lim_{s\to 1}\left(s-1\right)^{2}\zeta_{X\setminus\mathbb{P}^{2},-K_{X}}(s)=\lim_{s\to 1}\left(s-1\right)^{2}\zeta_{U,-K_{X}}(s),$
    hence the number of rational points of bounded anticanonical height in $X\setminus\mathbb{P}^{2}$ is “dominated” by the number of those points in the good open subset $U=U_{2}(a)$.
In both cases, we see that only after removing the closed subvariety $\mathbb{P}^{2}\subset X_{3}(a)$ we obtain a height zeta function $\zeta_{X\setminus\mathbb{P}^{2},-K_{X}}(s)$ satisfying the analogue of Conjecture 1.1 over global function fields.
We can also include here the easy case $a=0$ with $X=X_{3}(0)\simeq\mathbb{P}^{1}\times\mathbb{P}^{2}$. Indeed, it follows from Theorem 5.4 in Section 5.1 that $\zeta_{X,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$, has meromorphic continuation to $\mathbb{C}$ and has a pole of order 2 at $s=1$ with
$\lim_{s\to 1}(s-1)^{2}\zeta_{X,-K_{X}}(s)=\frac{q^{5(1-g)}h_{K}^{2}}{\zeta_{K}(3)\zeta_{K}(2)6(q-1)^{2}\log(q)^{2}}.$
In particular, in this easy case, there is no need to remove a closed subvariety of $X$ in order to verify the analogue of Conjecture 1.1 over global function fields.
Note that the above strategy also applies if we start with a Hirzebruch–Kleinschmidt variety $X$ with a height function induced by a general big metrized line bundle class $L\in\operatorname{Pic}(X)$. We refer the reader to Section 6 where this is illustrated in the case of Hirzebruch surfaces.
The techniques used in this paper are inspired by the work of Bourqui [Bou02] on the anticanonical height zeta function on Hirzebruch surfaces. Combining Bourqui’s ideas with some technical computations, we are able to express the height zeta function $\zeta_{U,L}(s)$, for the good open subset of a Hirzebruch–Kleinschmidt variety and for general big metrized line bundles $L$, essentially as a rational function of degree 2 in the zeta function $\zeta_{K}(s)$ of the base field $K$. In that regard, it would be interesting to investigate if this method can also be applied to other families of algebraic varieties defined over global function fields.
Remark 1.5.
The results obtained in this article served us as a guiding route for our investigations into the analogous problem for Hirzebruch–Kleinschmidt varieties over number fields, which are reported in the companion paper [HMM24].
This article is organized as follows. In Section 2 we introduce notation and recall some classical results concerning the geometry of curves defined over finite fields, and analytic properties the zeta functions of global function fields. We also include here a few general properties of toric varieties. In Section 3 we describe the effective cone, anticanonical class, and big divisor classes in Hirzebruch–Kleinschmidt varieties. In Section 4 we recall the analytic properties of the “standard height zeta” function of the projective space. The height zeta functions of Hirzebruch–Kleinschmidt varieties are studied in Section 5. In particular, our main theorems are presented in Section 5.1. Section 5 also includes a precise description of our decomposition of Hirzebruch–Kleinschmidt varieties (see Section 5.4) and the proof of Theorem 1.2 (see Section 5.5). Finally, in Section 6 we apply our general results to Hirzebruch surfaces. In particular, we recover the main result of [Bou02] as a particular case.
Acknowledgments
The authors thank Giancarlo Lucchini-Arteche for helpful comments on the first Galois cohomology group of an algebraic variety.
S. Herrero’s research is supported by ANID FONDECYT Iniciación grant 11220567 and by SNF grant CRSK-2_220746. T. Martinez’s research is supported by ANID FONDECYT Iniciación grant 11220567, ANID BECAS Doctorado Nacional 2020-21201321, and by UTFSM/DPP Programa de Incentivo a la Investigación Científica (PIIC) Convenio N°055/2023. P. Montero’s research is supported by ANID FONDECYT Regular grants 1231214 and 1240101.
2. Preliminaries
2.1. Basic notation
Let $\mathscr{C}$ be a projective, smooth, geometrically irreducible curve of genus $g$ defined over the finite field $\mathbb{F}_{q}$ of $q$ elements (as usual, $q$ is a positive power of a rational prime). Throughout this article we let $K$ denote the field $\mathbb{F}_{q}(\mathscr{C})$ of rational functions of $\mathscr{C}$. Associated to $K$ we have the following objects (see e.g. [Ros02] for details):
  • The set of valuations $\operatorname{Val}(K)$ of $K$, which is in bijection with the set of closed points of $\mathscr{C}$.
  • For each $v\in\operatorname{Val}(K)$, let $\mathscr{O}_{v}$ be the associated valuation ring and $m_{v}$ its maximal ideal, $k_{v}:=\mathscr{O}_{v}/m_{v}$ and $f_{v}:=[k_{v}:\mathbb{F}_{q}]$, so we have #$\#k_{v}=q^{f_{v}}$. We define for $x\in K$, $|x|_{v}:=q^{-f_{v}v(x)}$, so that for every $x\in K,x\neq 0$ we have the product formula
    $\prod_{v\in\operatorname{Val}(K)}|x|_{v}=1.$ (7)
  • The free abelian group $\operatorname{Div}(K)$ generated by $\operatorname{Val}(K)$. The elements of $\operatorname{Div}(K)$ are finite sums of the form $\sum_{v\in\operatorname{Val}(K)}n_{v}v$ with $n_{v}\in\mathbb{Z}$ and $n_{v}=0$ for all but finitely many $v\in\operatorname{Val}(K)$.
  • For $D=\sum_{v\in\operatorname{Val}(K)}n_{v}v\in\operatorname{Div}(K)$ and $v\in\operatorname{Val}(K)$ define $v(D):=n_{v}$ and set
    $\operatorname{Div}^{+}(K):=\{D\in\operatorname{Div}(K):v(D)\geq 0,\forall v\in\operatorname{Val}(K)\}.$
    Moreover, for $D_{1},\ldots,D_{n}\in\operatorname{Div}(K)$, we put
    $\sup(D_{1},\ldots,D_{n}):=\sum_{v\in\operatorname{Val}(K)}\sup\{v(D_{1}),\ldots,v(D_{n})\}v.$
  • For $x\in K^{\times}$, define the divisor of $x$ as $(x):=(x)_{0}-(x)_{\infty}$ where
    $(x)_{0}:=\sum_{\begin{subarray}{c}v\in\operatorname{Val}(K)\\ v(x)>0\end{subarray}}v(x)v,\quad(x)_{\infty}:=-\sum_{\begin{subarray}{c}v\in\operatorname{Val}(K)\\ v(x)<0\end{subarray}}v(x)v.$
    We also put $(0)=(0)_{0}=(0)_{\infty}=0$.
  • The degree function $\mathrm{\deg}:\operatorname{Div}(K)\to\mathbb{Z}$ defined by $\mathrm{\deg}(D):=\sum_{v\in\operatorname{Val}(K)}f_{v}v(D)$. We have $\deg((x))=0$ for all $x\in K$ (by the product formula (7) in the case $x\neq 0$).
  • The class number $h_{K}=\#(\operatorname{Div}^{0}(K)/(K))$, where $\operatorname{Div}^{0}(K):=\{D\in\operatorname{Div}(K):\deg(D)=0\}$ and $(K):=\{(x):x\in K\}$.
For a vector $\mathbf{a}=(a_{1},\ldots,a_{r})\in\mathbb{N}_{0}^{r}$, we write $|\mathbf{a}|:=\sum_{i=1}^{r}a_{i}$.
In this article all varieties will be assumed to be irreducible, reduced and separated schemes of finite type over the base field $K$. For simplicity we write $\mathbb{A}^{n}$ and $\mathbb{P}^{n}$ for the affine and projective space of dimension $n$ over $K$, and products of varieties are to be understood as fiber products over $\mathrm{Spec}(K)$.
2.2. Geometric and arithmetic tools
For $D\in\operatorname{Div}(K)$, let $\ell(D):=\dim_{\mathbb{F}_{q}}H^{0}(\mathscr{C},\mathscr{O}_{\mathscr{C}}(D))$, where
$H^{0}(\mathscr{C},\mathscr{O}_{\mathscr{C}}(D)):=\{x\in K^{\times}:(x)+D\geq 0\}\cup\{0\}.$
Note that $\ell(D)=0$ if $\deg(D)<0$ and if $D\geq 0$ then #$\#\{x\in K:(x)_{\infty}\leq D\}=q^{\ell(D)}$.
As a consequence of the Riemann–Roch theorem for curves (see e.g. [Ros02, Theorem 5.4]) we have that
$\ell(D)-\ell(K_{\mathscr{C}}-D)=\deg(D)+1-g,$
where $K_{\mathscr{C}}$ is a canonical divisor class of the curve $\mathscr{C}$. In particular, if $\deg(D)>2g-2=\deg(K_{\mathscr{C}})$ then
$\ell(D)=\deg(D)+1-g.$
We also define the zeta function
$\operatorname{Z}_{K}(T):=\sum_{D\geq 0}T^{\deg(D)}=\prod_{v\in\operatorname{Val}(K)}\left(1-T^{f_{v}}\right),$
and define the zeta function of $K$ as
$\zeta_{K}(s):=\operatorname{Z}_{K}(q^{-s}).$
The series defining $\zeta_{K}(s)$ converges absolutely and uniformly on compact subsets of the half-plane $\Re(s)>1$ and has no zeroes in that domain. Moreover, it is a rational function in $q^{-s}$ and has a simple pole at $s=1$ with residue
$\operatorname{Res}_{s=1}\zeta_{K}(s)=\frac{h_{K}q^{1-g}}{(q-1)\log(q)}$
(see e.g. [Ros02, Theorem 5.9]).
2.3. Toric varieties
We refer the reader to [CLS11] for the general theory of toric varieties. Here, we simply recall that a toric variety is a normal and separated variety that contains a maximal torus $\mathbb{T}$ as an open subset and admits an effective regular action of $\mathbb{T}$ extending the natural action of this torus over itself. The relevant toric varieties appearing in this paper are all smooth. We recall that on a smooth variety $X$ every Weil divisor is Cartier, and in particular $\operatorname{Cl}(X)\simeq\operatorname{Pic}(X)$.
3. Hirzebruch–Kleinschmidt varieties
Given integers $r\geq 1,t\geq 2,0\leq a_{1}\leq\ldots\leq a_{r},$ consider the vector bundle
$\mathscr{E}:=\mathscr{O}_{\mathbb{P}^{t-1}}\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(a_{1})\oplus\cdots\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(a_{r}).$
Associated to $\mathscr{E}$ there is a projective bundle $\pi:\mathbb{P}(\mathscr{E})\to\mathbb{P}^{t-1}$. The variety $\mathbb{P}(\mathscr{E})$ is smooth, projective and toric, of dimension
$d:=\dim(\mathbb{P}(\mathscr{E}))=r+t-1,$
and has $\operatorname{Pic}(\mathbb{P}(\mathscr{E}))\simeq\mathbb{Z}^{2}$ (see [CLS11, Chapter 7] or [HMM24, Section 3]). Conversely, Kleinschmidt [Kle88] proved the following classification result of smooth projective [^2] Actually, the classification in [Kle88] does not assume the projectivity hypothesis. toric varieties of Picard rank 2, and gave explicit defining equations in [Kle88, Theorem 3].
Theorem 3.1 (Kleinschmidt).
Let $X$ be a smooth projective toric variety with $\operatorname{Pic}(X)\simeq\mathbb{Z}^{2}$. Then, there exists integers $r\geq 1,t\geq 2,0\leq a_{1}\leq\cdots\leq a_{r}$ with $r+t-1=d=\dim(X)$ such that
$X\simeq\mathbb{P}(\mathscr{O}_{\mathbb{P}^{t-1}}\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(a_{1})\oplus\cdots\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(a_{r})).$
Moreover, if we write $I_{k}:=\{1,\ldots,k\}\subset\mathbb{Z}$, then $X$ is isomorphic to the subvariety $X_{d}(a_{1},\ldots,a_{r})\subset\mathbb{P}^{rt}\times\mathbb{P}^{t-1}$ given in homogeneous coordinates $([x_{0}:(x_{ij})_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t}])$ by the equations
$x_{mj}y^{a_{j}}_{n}=x_{nj}y^{a_{j}}_{m},\,\text{ for all }j\in I_{r}\text{ andall }m,n\in I_{t}\text{ with }m\neq n.$ (8)
Recall from the introduction that we define Hirzebruch–Kleinschmidt varieties exactly as the family of algebraic varieties $X_{d}(a_{1},\ldots,a_{r})$ appearing in Theorem 3.1.
3.1. Effective divisors
In this section we give a basis of $\operatorname{Pic}(X)$ for a Hirzebruch–Kleinschmidt variety $X$, and use it to describe the anticanonical class, the effective cone, and the big divisor classes.
Proposition 3.2.
Let $X=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety and let us denote by $f$ the class of $\pi^{\ast}\mathscr{O}_{\mathbb{P}^{t-1}}(1)$ and by $h$ the class of $\mathscr{O}_{X}(1)$, both in $\operatorname{Pic}(X)$. Then:
  1. (1)
    $\operatorname{Pic}(X)=\mathbb{Z}h\oplus\mathbb{Z}f$.
  2. (2)
    The anticanonical divisor class of $X$ is given by
    $-K_{X}=(r+1)h+\left(t-|\mathbf{a}|\right)f,$
    where $|\mathbf{a}|=\sum_{i=1}^{r}a_{i}$.
  3. (3)
    The cone of effective divisors of $X$ is given by
    $\Lambda_{\textup{eff}}(X)=\{\gamma h+\xi f:\gamma\geq 0,\xi\geq-\gamma a_{r}\}\subset\operatorname{Pic}(X)_{\mathbb{R}}$
    where $\operatorname{Pic}(X)_{\mathbb{R}}:=\operatorname{Pic}(X)\otimes_{\mathbb{Z}}\mathbb{R}$.
Proof.
The analogue statement for the variety
$Y:=\mathbb{P}(\mathscr{O}_{\mathbb{P}^{t-1}}\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(-a_{r})\oplus\cdots\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(a_{r-1}-a_{r}))$
was given in [HMM24, Proposition 3.4]. More precisely, putting $f_{Y}:=\pi_{Y}^{\ast}\mathscr{O}_{\mathbb{P}^{t-1}}(1)$ and $h_{Y}:=\mathscr{O}_{Y}(1)$, where $\pi_{Y}:Y\to\mathbb{P}^{t-1}$ is the natural projection, it was proven that
$\operatorname{Pic}(Y)$ $=$ $\mathbb{Z}h_{Y}\oplus\mathbb{Z}f_{Y},$
$-K_{Y}$ $=$ $(r+1)h_{Y}+\left((r+1)a_{r}+t-|\mathbf{a}|\right)f_{Y},\text{ and}$
$\Lambda_{\textup{eff}}(Y)$ $=$ $\{\gamma h_{Y}+\xi f_{Y}:\gamma\geq 0,\xi\geq 0\}.$
Now, under the canonical isomorphism of projective bundles
$Y\simeq\mathbb{P}((\mathscr{O}_{\mathbb{P}^{t-1}}\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(-a_{r})\oplus\cdots\oplus\mathscr{O}_{\mathbb{P}^{t-1}}(a_{r-1}-a_{r}))\otimes\mathscr{O}_{\mathbb{P}^{t-1}}(a_{r}))=X_{d}(a_{1},\ldots,a_{r}),$
the line bundle classes $-K_{Y},f_{Y}$ and $h_{Y}$ correspond to $-K_{X},f$ and $h-a_{r}f$, respectively (see e.g. [Har77, Chapter V, Corollary 2.18]). Then, simple computations lead to the desired results. We omit the details for brevity. ∎
It follows from [Laz04, Theorem 2.2.26] that a divisor class in $\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$ is big if and only if it lies in the interior of the effective cone $\Lambda_{\textup{eff}}(X)$. Hence, we get the following corollary from Proposition 3.2.
Corollary 3.3.
Let $L=\gamma h+\xi f$ with $\gamma,\xi\in\mathbb{Z}$, where $\{h,f\}$ is the basis of $\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$ given in Proposition 3.2. Then, $L$ is big if and only if $\gamma>0$ and $\xi>-\gamma a_{r}$.
3.2. Peyre’s $\alpha^{\ast}$-constant
Let us recall that Peyre’s $\alpha^{\ast}$-constant of an almost Fano [^3] Every smooth projective toric variety is almost Fano. variety $X$ over a global function field can be defined as
$\alpha^{\ast}(X)=\int_{\Lambda_{\textup{eff}}(X)^{\vee}}e^{-\left\langle-K_{X},\mathbf{y}\right\rangle}\textup{d}\mathbf{y},$
[…]
Lemma 3.4.
Let $X=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety. Then, its $\alpha^{\ast}$-constant is given by
$\alpha^{\ast}(X)=\int_{0}^{\infty}\int_{a_{r}y_{2}}^{\infty}e^{-(r+1)y_{1}-(t-|\mathbf{a}|)y_{2}}\textup{d}y_{1}\,\textup{d}y_{2}=\frac{1}{(r+1)\left((r+1)a_{r}+t-|\mathbf{a}|\right)}.$
4. Height zeta function of the projective space
For an integer $n\geq 1$, we consider in this section the projective space $\mathbb{P}^{n}$ over $K$ with standard height function
$H_{\mathbb{P}^{n}}([x_{0},\ldots,x_{n}]):=\prod_{v\in\operatorname{Val}(K)}\sup\{|x_{0}|_{v},|x_{1}|_{v},\ldots,|x_{n}|_{v}\},$ (9)
and standard height zeta function
$\zeta_{\mathbb{P}^{n}}(s):=\sum_{P\in\mathbb{P}^{n}(K)}H_{\mathbb{P}^{n}}(P)^{-s}\quad\text{for }s\in\mathbb{C}\text{ with }\Re(s)\gg 0.$
The goal of this section is to state the following well-known theorem, for which we give a proof for convenience of the reader.
Theorem 4.1.
The height zeta function $\zeta_{\mathbb{P}^{n}}(s)$ converges absolutely for $\Re(s)>n+1$ and it is a rational function in $q^{-s}$. Moreover, it has a simple pole at $s=n+1$ with residue
$\operatorname{Res}_{s=n+1}\zeta_{\mathbb{P}^{n}}(s)=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)\log(q)}.$
Proof.
The fact that $\zeta_{\mathbb{P}^{n}}(s)$ is a rational function in $q^{-s}$ follows from [Wan92, Theorem 3.2]. Now, for $m\geq 1$ put $N(m):=\#\{P\in\mathbb{P}^{n}(K):H(P)=m\}$. By definition of the standard height function, we have $N(m)=0$ if $m$ is not of the form $m=q^{d}$ with $d\geq 0$ an integer. Hence
$\zeta_{\mathbb{P}^{n}}(s)=\sum_{d\geq 0}N(q^{d})q^{-ds}.$
Now, by [Ser89, Section 2.5] we have the estimate
$N(q^{d})=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)}q^{d(n+1)}+O(q^{dn})\quad\text{as }d\to\infty.$ (10)
This implies that $\zeta_{\mathbb{P}^{n}}(s)$ converges absolutely for $\Re(s)>n+1$, and moreover
$\zeta_{\mathbb{P}^{n}}(s)=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)}\left(\frac{1}{1-q^{(n+1)-s}}\right)+F(s),$
with $F(s)$ an analytic function on $\Re(s)>n$. Hence, we also get
$\operatorname{Res}_{s=n+1}\zeta_{\mathbb{P}^{n}}(s)=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)}\left(\lim_{s\to n+1}\frac{s-(n+1)}{1-q^{(n+1)-s}}\right)=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)\log(q)}.$
This shows the desired results and completes the proof of the theorem. ∎
Remark 4.2.
  1. (1)
    A better estimate for the error term in (10) is given in [Wan92, Corollary 4.3], which is related to the fact that $\zeta_{\mathbb{P}^{n}}(s)$ is analytic for $\Re(s)\in\left]\frac{1}{2},n+1\right[$. For the purpose of this paper, however, we do not require such a strong result.
  2. (2)
    Theorem 4.1 also follows from [Pey12, Théorème 3.11].
5. Height zeta functions of Hirzebruch–Kleinschmidt varieties
In this section, we use the explicit equations defining our Hirzebruch–Kleinschmidt varieties in order to study the analytic properties of their height zeta functions. To do this, we first present some lemmas concerning Möbius inversion, supremum of divisors and relations between different zeta functions. We will first focus on the case $a_{r}>0$ in the family of varieties $X=X_{d}(a_{1},\ldots,a_{r})$, and later treat the much easier case where $a_{r}=0$, since then $X\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ and the relevant height zeta functions are just products of height zeta functions of projective spaces.
In the notation of Proposition 3.2, let $L=\gamma h+\xi f$ be the class of a big line bundle. Then, the height function induced by $L$ on the set of rational points of $X_{d}(a_{1},\ldots,a_{r})$ is
$H_{L}\left(([x_{0}:(x_{ij})_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t}])\right):=\prod_{v\in\operatorname{Val}(K)}\sup_{i\in I_{t},j\in I_{r}}\left\{|x_{0}|_{v},|x_{ij}|_{v}\right\}^{\gamma}\prod_{v\in\operatorname{Val}(K)}\sup_{i\in I_{t}}\left\{|y_{i}|_{v}\right\}^{\xi}.$ (11)
It is easy to see that this height function is associated to an adelic metrization of $L$ in the sense of [Pey12, Secion 1.2]. We refer to this as the “standard metrization” of $L$.
As explained in the introduction, to carry out the analysis of the corresponding height zeta functions, we find it necessary to remove a closed subset in order to avoid accumulation of rational points of bounded height. To this purpose, the following definition was given in the introduction.
Definition 5.1.
Given integers $r\geq 1$, $t\geq 2$ and $0\leq a_{1}\leq\cdots\leq a_{r}$ defining the Hirzebruch–Kleinschmidt variety $X_{d}(a_{1},\ldots,a_{r})$ as in Theorem 3.1, we define the good open subset $U_{d}(a_{1},\ldots,a_{r})\subset X_{d}(a_{1},\ldots,a_{d})$ as the complement of the closed subvariety defined by the equation $x_{tr}=0$.
5.1. Height zeta functions
Let $L=\gamma h+\xi f\in\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$ big. The height zeta function associated to $H_{L}$ over the good open subset $U:=U_{d}(a_{1},\ldots,a_{r})$ is defined as
$\zeta_{U,L}(s):=\sum_{x\in U(K)}H_{L}(x)^{-s}.$
Associated to $L$ as above, we also define
$A_{L}:=\frac{r+1}{\gamma},\quad B_{L}:=\frac{(r+1)a_{r}-|\mathbf{a}|+t}{\gamma a_{r}+\xi}.$ (12)
Note that by Proposition 3.2 we have
$a(L)=\max\{A_{L},B_{L}\}\quad\text{and}\quad b(L)=\left\{\begin{array}[]{ll}2&\text{if }A_{L}=B_{L},\\ 1&\text{if }A_{L}\neq B_{L}.\end{array}\right.$ (13)
We now present the first main result of this section, which describes the analytical properties of the function $\zeta_{U,L}(s)$ in the case where $a_{r}>0$.
Theorem 5.2.
Let $X:=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$ with $a_{r}>0$. Moreover, let $U:=U_{d}(a_{1},\ldots,a_{r})$ be the good open subset of $X$, let $L=\gamma h+\xi f$ be a big line bundle class in $\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$, and let $\zeta_{U,L}(s)$ be the associated height zeta function. Then, $\zeta_{U,L}(s)$ is a rational function in $q^{-s}$. Moreover, $\zeta_{U,L}(s)$ converges absolutely for $\Re(s)>a(L)$ and it has a pole of order $b(L)$ at $s=a(L)$ with
$\lim_{s\to a(L)}(s-a(L))^{b(L)}\zeta_{U,L}(s)=\left\{\begin{array}[]{ll}\vspace{0.2cm}\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(\gamma a_{r}+\xi)\gamma(q-1)^{2}\log(q)^{2}}&\text{if }\xi=\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma,\\ \vspace{0.2cm}\frac{q^{(d+2-N_{X})(1-g)}h_{K}\mathbf{R}_{K}(1-N_{X},\gamma B_{L}-r+N_{X}-1)}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma B_{L})(\gamma a_{r}+\xi)(q-1)\log(q)}&\text{if }\xi<\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma,\\ \vspace{0.2cm}\frac{q^{(r+1)(1-g)}h_{K}\mathbf{R}_{K}\left(1-t,A_{L}\xi+|\mathbf{a}|\right)}{\zeta_{K}\left(A_{L}\xi+|\mathbf{a}|\right)\zeta_{K}(r+1)\gamma(q-1)\log(q)}&\text{if }\xi>\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma,\\ \end{array}\right.$
where $N_{X}:=\#\{j\in\{1,\ldots,r\}:a_{j}=a_{r}\}$ and
$\mathbf{R}_{K}(a,b):=\sum_{D\geq 0}q^{-(a\ell(D)+b\deg(D))}\quad\text{for }a,b\in\mathbb{C}\text{ with }\Re(a+b)>1.$
The proof of Theorem 5.2 is given in Section 5.3, after establishing some preliminary lemmas in Section 5.2.
Remark 5.3.
  1. (1)
    The Riemann–Roch theorem (see Section 2.2) implies that
    $\mathbf{R}_{K}(a,b)=\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{-(a\ell(D)+b\deg(D))}+q^{a(g-1)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-(a+b)\deg(D)},$ (14)
    hence $\mathbf{R}_{K}(a,b)=q^{a(g-1)}\zeta_{K}(a+b)+\mathbf{S}_{K}(a,b)$ with the finite sum
    $\mathbf{S}_{K}(a,b):=\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}(q^{-(a\ell(D)+b\deg(D))}-q^{a(g-1)-(a+b)\deg(D)}).$
    In particular, $\mathbf{R}_{K}(a,b)=q^{-a}\zeta_{K}(a+b)$ if $g=0$.
  2. (2)
    As in the case of Hirzebruch–Kleinschmidt varieties over number fields (see [HMM24, Remark 6.8]), the different cases that appear in Theorem 5.2 give a subdivision of the big cone of $X$, i.e. the interior of $\Lambda_{\textup{eff}}(X)$ (see Figure 1 for an illustration, where we assume $t>|\mathbf{a}|$ for simplicity). It follows from the proof of Theorem 5.2 in Section 5.3 that the line bundles $L$ contained in the ray passing through the anticanonical class have height zeta functions with a double pole at $s=A_{L}=B_{L}$, while line bundles outside this ray have height zeta functions with a simple pole at $s=\max\{A_{L},B_{L}\}$.
$\xi$$\gamma$$-K_{X}$$\xi=\frac{t-|\mathbf{a}|}{r+1}\gamma$$\xi<\frac{t-|\mathbf{a}|}{r+1}\gamma$$\xi>\frac{t-|\mathbf{a}|}{r+1}\gamma$
Figure 1. Subdivision of the big cone of $X_{d}(a_{1},\ldots,a_{r})$ (when $t>|\mathbf{a}|$ for simplicity).
The much simpler case when $a_{r}=0$ is given by the following theorem, where we see that there is no need to remove a closed subvariety of $X=X_{d}(a_{1},\ldots,a_{r})$ and we can directly give the analytic properties of the height zeta function
$\zeta_{X,L}(s):=\sum_{P\in X(K)}H_{L}(P)^{-s}.$
Note that, in this case, we have
$A_{L}=\frac{r+1}{\gamma},\quad B_{L}=\frac{t}{\xi}.$
Theorem 5.4.
Let $X\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ be a Hirzebruch–Kleinschmidt variety over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$ with $a_{r}=0$ and let $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big. Then, the height zeta function $\zeta_{X,L}(s)$ is a rational function in $q^{-s}$. Moreover, $\zeta_{X,L}(s)$ converges absolutely for $\Re(s)>a(L)$ and it has a pole of order $b(L)$ at $s=a(L)$ with
$\lim_{s\to a(L)}(s-a(L))^{b(L)}\zeta_{X,L}(s)=\left\{\begin{array}[]{ll}\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)\gamma\xi(q-1)^{2}\log(q)^{2}}&\text{if }\xi=\left(\frac{t}{r+1}\right)\gamma,\vspace{0.1cm}\\ \frac{\zeta_{\mathbb{P}^{r}}(\gamma B_{L})h_{K}q^{t(1-g)}}{\zeta_{K}(t)\xi(q-1)\log(q)}&\text{if }\xi<\left(\frac{t}{r+1}\right)\gamma,\vspace{0.1cm}\\ \frac{\zeta_{\mathbb{P}^{t-1}}(\xi A_{L})h_{K}q^{(r+1)(1-g)}}{\gamma\zeta_{K}(r+1)(q-1)\log(q)}&\text{if }\xi>\left(\frac{t}{r+1}\right)\gamma.\vspace{0.1cm}\end{array}\right.$
Proof.
[…]
$\zeta_{X,L}(s)=\zeta_{\mathbb{P}^{r}}(\gamma s)\zeta_{\mathbb{P}^{t-1}}(\xi s).$
Then, using Theorem 4.1, we see that $\zeta_{X,L}(s)$ is a rational function in $q^{-s}$ and it converges absolutely for $\Re(s)>\max\{A_{L},B_{L}\}=a(L)$. Moreover, we have the following properties:
  1. (1)
    If $A_{L}=B_{L}$, then $\zeta_{X,L}(s)$ has a double pole at $s=a(L)$ with
    $\begin{split}\lim_{s\to a(L)}(s-a(L))^{2}\zeta_{X,L}(s)&=\frac{\operatorname{Res}_{s=r+1}\zeta_{\mathbb{P}^{r}}(s)\operatorname{Res}_{s=t}\zeta_{\mathbb{P}^{t-1}}(s)}{\gamma\xi}\\ &=\frac{h_{K}^{2}q^{(d+2)(1-g)}}{\gamma\xi\zeta_{K}(r+1)\zeta_{K}(t)(q-1)^{2}\log(q)^{2}}.\end{split}$
  2. (2)
    If $A_{L}<B_{L}$, then $\zeta_{X,L}(s)$ has a simple pole at $s=B_{L}=a(L)$ with
    $\begin{split}\lim_{s\to a(L)}(s-a(L))\zeta_{X,L}(s)&=\frac{\zeta_{\mathbb{P}^{r}}(\gamma B_{L})\operatorname{Res}_{s=t}\zeta_{\mathbb{P}^{t-1}}(s)}{\xi}\\ &=\frac{\zeta_{\mathbb{P}^{r}}(\gamma B_{L})h_{K}q^{t(1-g)}}{\xi\zeta_{K}(t)(q-1)\log(q)}.\end{split}$
  3. (3)
    If $A_{L}>B_{L}$, then $\zeta_{X,L}(s)$ has a simple pole at $s=A_{L}=a(L)$ with
    $\begin{split}\lim_{s\to a(L)}(s-a(L))\zeta_{X,L}(s)&=\frac{\operatorname{Res}_{s=r+1}\zeta_{\mathbb{P}^{r}}(s)\zeta_{\mathbb{P}^{t-1}}(\xi A_{L})}{\gamma}\\ &=\frac{\zeta_{\mathbb{P}^{t-1}}(\xi A_{L})h_{K}q^{(r+1)(1-g)}}{\gamma\zeta_{K}(r+1)(q-1)\log(q)}.\end{split}$
These cases correspond exactly to $\xi$ equal, less than, or greater than $\left(\frac{t}{r+1}\right)\gamma$, respectively. This proves the theorem. ∎
5.2. Preliminary lemmas
In this section we stablish some preliminary lemmas needed for the proof of Theorem 5.2 in the next section.
Definition 5.5.
Given functions $f,g:\operatorname{Div}^{+}(K)\rightarrow\mathbb{C}$, we define their convolution product as
$(f\star g)(D):=\sum_{0\leq D^{\prime}\leq D}f(D^{\prime})g(D-D^{\prime}).$
Moreover, we define the functions $1,\delta,\mu:\operatorname{Div}^{+}(K)\rightarrow\mathbb{C}$ by
  • $11(D):=$ 1 for every $D\in\operatorname{Div}^{+}(K)$.
  • $\delta(D):=\left\{\begin{array}[]{ll}1&\textup{if }D=0,\\ 0&\textup{otherwise}.\end{array}\right.$
  • $\mu(D):=\left\{\begin{array}[]{ll}1&\textup{if }D=0,\\ (-1)^{\sum_{v}v(D)}&\textup{if $D\neq 0$ and $v(D)=0$ or $1$ for all $v\in\operatorname{Val}(K)$},\\ 0&\textup{otherwise}.\end{array}\right.$
In particular, $\delta$ is a unit for the convolution product and $1\star\mu=\mu\star 1=\delta$.
The following result, based on a straightforward computation, is a Möbius inversion formula in this context.
Lemma 5.6.
Let $D\in\operatorname{Div}^{+}(K)$ and let $f,g:\operatorname{Div}^{+}(K)\to\mathbb{C}$ be two functions. Then, the relation
$f(D)=\sum_{0\leq D^{\prime}\leq D}g(D^{\prime})$
is equivalent to
$g(D)=\sum_{0\leq D^{\prime}\leq D}\mu(D-D^{\prime})f(D^{\prime}).$
Lemma 5.6 motivates the following definition that will be useful for counting purposes.
Definition 5.7.
Let $f,g:\operatorname{Div}^{+}(K)\to\mathbb{C}$ be functions. We say that the pair $(f,g)$ forms a $\mu$-couple if they satisfy any of the two equivalent relations in Lemma 5.6.
By definition of the function $\delta$, the pair $(\delta,1)$ forms a $\mu$-couple.
Lemma 5.8.
We have the following equality of formal series:
$\sum_{D\geq 0}\mu(D)T^{\deg(D)}=\frac{1}{\operatorname{Z}_{K}(T)}.$
Proof.
Since $(\delta,1)$ is a $\mu$-couple, we have that
$\begin{split}\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D^{\prime})}\operatorname{Z}_{K}(T)&=\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D^{\prime})}\sum_{D\geq 0}T^{\deg(D)}\\ &=\sum_{D\geq 0}\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D+D^{\prime})}\\ &=\sum_{D\geq 0}\sum_{0\leq D^{\prime}\leq D}\mu(D-D^{\prime})T^{\deg(D)}\\ &=\sum_{D\geq 0}\delta(D)T^{\deg(D)}=1,\end{split}$
where the last equality follows from the definition of the function $\delta$. This proves the lemma. ∎
Putting $T=q^{-s}$ in the above lemma, we get the following corollary.
Corollary 5.9.
For every $s\in\mathbb{C}$ with $\Re(s)>1$, we have
$\sum_{D\geq 0}\mu(D)q^{-s\deg(D)}=\frac{1}{\zeta_{K}(s)},$
and the series on the left hand side converges absolutely and uniformly on compact subsets of the half-plane $\Re(s)>1$.
For computations involving supremum of divisors, we will make use of the following result.
Lemma 5.10.
For all $x,y\in K$ we have
$\sup\{(x)_{\infty},(xy)_{\infty},(y)_{\infty}\}=(x)_{\infty}+(y)_{\infty}.$
Proof.
Let $v\in\operatorname{Val}(K)$. We recall that for a divisor $D=\sum_{v\in\operatorname{Val}(K)}n_{v}v\in\operatorname{Div}(K)$, we defined $v(D)=n_{v}$. We have to show that
$\sup\{v((x)_{\infty}),v((xy)_{\infty}),v((y)_{\infty})\}=v((x)_{\infty})+v((y)_{\infty}),$
for all $v\in\operatorname{Val}(K)$. We have the following four cases to consider:
  1. (1)
    If $v(x),v(y)\leq 0$ then $v\left((xy)_{\infty}\right)=v\left((x)_{\infty}\right)+v\left((y)_{\infty}\right)$.
  2. (2)
    If $v(x)\leq 0$ and $v(y)>0$ then $v\left((xy)_{\infty}\right)\leq v\left((x)_{\infty}\right)$ and $v\left((y)_{\infty}\right)=0$.
  3. (3)
    If $v(x)>0$ and $v(y)\leq 0$ then $v\left((xy)_{\infty}\right)\leq v\left((y)_{\infty}\right)$ and $v\left((x)_{\infty}\right)=0$.
  4. (4)
    If $v(x),v(y)>0$ then $v\left((xy)_{\infty}\right)=v\left((x)_{\infty}\right)=v\left((y)_{\infty}\right)=0$.
The desired results follows directly from these computations. This proves the lemma. ∎
Lemma 5.11.
The good open subset $U_{d}(a_{1},\ldots,a_{r})$ equals the image of the map
$(x_{1},\ldots,x_{d})\in\mathbb{A}^{d}\mapsto([x_{d}:(\tilde{x}_{ij})_{i\in I_{t},j\in I_{r}}],[x_{1}:\ldots:x_{t-1}:1])\in X_{d}(a_{1},\ldots,a_{r}),$ (15)
where
$\tilde{x}_{ij}=\left\{\begin{array}[]{ll}x_{t+j-1}x_{i}^{a_{j}}&\text{if }i<t,j<r,\\ x_{t+j-1}&\text{if }i=t,j<r,\\ x_{i}^{a_{r}}&\text{if }i<t,j=r,\\ 1&\text{if }i=t,j=r.\end{array}\right.$
Moreover, the map (15) is injective.
Proof.
Let $P:=([x_{0}:(x_{ij})_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t}])$ be a point in $U_{d}(a_{1},\ldots,a_{r})$, and note that the condition $x_{tr}\neq 0$ implies $y_{t}\neq 0$. Indeed, due to the equations
$x_{tr}y_{n}^{a_{r}}=x_{nr}y_{t}^{a_{r}}\quad\text{for }n=1,\ldots,t-1,$
the vanishing of $y_{t}$ would imply that all the coordinates of $[y_{1}:\ldots:y_{t}]\in\mathbb{P}^{t-1}$ would be zero, which is absurd. Hence, we can assume $y_{t}=x_{tr}=1$. Putting $x_{i}:=y_{i}$ for $1\leq i<t$, $x_{t+j-1}:=x_{tj}$ for $1\leq j<r$, and $x_{d}:=x_{0}$, we obtain a point $(x_{1},\ldots,x_{d})\in\mathbb{A}^{d}$ which defines coefficients $\tilde{x}_{ij}$ as above. By construction, we have $\tilde{x}_{ij}=x_{ij}$ for $i=t$ and for all $j\in I_{r}$. Now, it follows from the equations (8) with $n=t$ and $j<r$ that $x_{ij}=x_{tj}y_{i}^{a_{j}}=x_{t+j-1}x_{i}^{a_{j}}=\tilde{x}_{ij}$ for all $i$ with $1\leq i<t$. Similarly, using the equations (8) with $n=t$ and $j=r$, we get $x_{ir}=x_{tr}y_{i}^{a_{r}}=x_{i}^{a_{r}}=\tilde{x}_{ir}$ for all $i$ with $1\leq i<t$. This shows that the image of $(x_{1},\ldots,x_{d})$ under (15) equals $P$, and proves the first statement of the lemma. Finally, the injectivity of the map (15) follows from the fact that one can recover $(x_{1},\ldots,x_{d})$ from the coefficients $x_{0}$, $x_{tj}$ with $1\leq j<r$, and $y_{k}$ for $1\leq k<t$, of $P$. This completes the proof of the lemma. ∎
We now introduce the formal power series
$\operatorname{Z}_{U,L}(T):=\sum_{x\in\mathbb{A}^{d}}T^{\deg(\mathbf{d}_{L}(x))},$ (16)
where for $x=(x_{i})\in\mathbb{A}^{d}$ we put
$\mathbf{d}_{L}(x):=\gamma\sup_{\begin{subarray}{c}i\in I_{t-1}\\ j\in I_{r-1}\end{subarray}}\big{(}(x_{d})_{\infty},(x_{t-1+j})_{\infty}+a_{j}(x_{i})_{\infty},a_{r}(x_{i})_{\infty}\big{)}+\xi\sup_{i\in I_{t-1}}\big{(}(x_{i})_{\infty}\big{)}.$
As in the case of the zeta function of the base field $K$ (see Section 2.2), we have the following relation.
Lemma 5.12.
With the above notation, we have the equality
$\zeta_{U,L}(s)=\operatorname{Z}_{U,L}(q^{-s}).$
Proof.
Let $x=(x_{i})\in\mathbb{A}^{d}$ and denote by $P_{x}$ the image of $x$ in $U$ under the map defined in Lemma 5.11. Then, we have
$H_{L}(P_{x})$ $=\prod_{v}\sup_{\begin{subarray}{c}i\in I_{t-1}\\ j\in I_{r-1}\end{subarray}}\left\{|x_{d}|_{v},|x_{t-1+j}x_{i}^{a_{j}}|_{v},|x_{t-1+j}|_{v},|x_{i}^{a_{r}}|_{v},1\right\}^{\gamma}\prod_{v}\sup_{i\in I_{t-1}}\left\{|x_{i}|_{v},1\right\}^{\xi},$
where each product runs over all $v\in\operatorname{Val}(K)$. Putting
$V_{1}$ ::=:= $\left\{v\in\operatorname{Val}(K):\inf_{\begin{subarray}{c}i\in I_{t-1}\\ j\in I_{r-1}\end{subarray}}\left\{v(x_{d}),v(x_{t-1+j}x_{i}^{a_{j}}),v(x_{t-1+j}),v(x_{i}^{a_{r}})\right\}<0\right\},$
$V_{2}$ ::=:= $\left\{v\in\operatorname{Val}(K):\inf_{i\in I_{t-1}}\left\{v(x_{i})\right\}<0\right\},$
we get
$H_{L}(P_{x})$ $=\prod_{v\in V_{1}}q^{-f_{v}\gamma\inf_{\begin{subarray}{c}i\in I_{t-1}\\ j\in I_{r-1}\end{subarray}}\left\{v(x_{d}),v(x_{t-1+j}x_{i}^{a_{j}}),v(x_{t-1+j}),v(x_{i}^{a_{r}})\right\}}\prod_{v\in V_{2}}q^{-f_{v}\xi\inf_{i\in I_{i-1}}\left\{v(x_{i})\right\}}.$
By taking $\log_{q}$ on both sides, we obtain
$\log_{q}H_{L}(P_{x})$ $=\gamma\sum_{v\in V_{1}}-f_{v}\inf_{\begin{subarray}{c}i\in I_{t-1}\\ j\in I_{r-1}\end{subarray}}\left\{v(x_{d}),v(x_{t-1+j}x_{i}^{a_{j}}),v(x_{t-1+j}),v(x_{i}^{a_{r}})\right\}$
$\hskip 85.35826pt+\xi\sum_{v\in V_{2}}-f_{v}\inf_{i\in I_{t-1}}\left\{v(x_{i})\right\}$
$=\gamma\deg\left(\sup_{\begin{subarray}{c}i\in I_{t-1}\\ j\in I_{r-1}\end{subarray}}\big{(}(x_{d})_{\infty},(x_{t-1+j}x_{i}^{a_{j}})_{\infty},(x_{t-1+j})_{\infty},a_{r}(x_{i})_{\infty}\big{)}\right)$
$\hskip 85.35826pt+\xi\deg\left(\sup_{i\in I_{t-1}}\big{(}(x_{i})_{\infty}\big{)}\right).$
Since $a_{j}(x_{i})_{\infty}\leq a_{r}(x_{i})_{\infty}$, we can add $a_{j}(x_{i})_{\infty}$ inside the first supremum and use Lemma 5.10 to get
$\log_{q}H_{L}(P_{x})$ $=\gamma\deg\left(\sup_{\begin{subarray}{c}i\in I_{t-1}\\ j\in I_{r-1}\end{subarray}}\big{(}(x_{d})_{\infty},(x_{t-1+j})_{\infty}+a_{j}(x_{i})_{\infty},a_{r}(x_{i})_{\infty}\big{)}\right)$
$\hskip 85.35826pt+\xi\deg\left(\sup_{i\in I_{t-1}}\big{(}(x_{i})_{\infty}\big{)}\right)$
$=\deg(\mathbf{d}_{L}(x)).$
It follows that $H_{L}(P_{x})^{-s}=q^{-s\deg(\mathbf{d}_{L}(x))}$, hence
$\zeta_{U,L}(s)=\sum_{x\in\mathbb{A}^{d}}H_{L}(P_{x})^{-s}=\sum_{x\in\mathbb{A}^{d}}q^{-s\deg(\mathbf{d}_{L}(x))}=\operatorname{Z}_{U,L}(q^{-s})$
by Lemma 5.11. This completes the proof of the lemma. ∎
Analogous to Bourqui in [Bou02, p. 351], we define the counting functions
$\begin{split}\tilde{R}_{L}(D):=&\#\left\{x\in\mathbb{A}^{d}:\mathbf{d}_{L}(x)=D\right\},\\ R_{L}(D):=&\#\left\{x\in\mathbb{A}^{d}:\mathbf{d}_{L}(x)\leq D\right\}.\end{split}$
Remark 5.13.
Note that $R_{L}(D)=\sum_{0\leq D^{\prime}\leq D}\tilde{R}_{L}(D^{\prime})$, so $(R_{L},\tilde{R}_{L})$ forms a $\mu$-couple.
It follows from the definition of $\operatorname{Z}_{U,L}(T)$ in (16) that
$\operatorname{Z}_{U,L}(T)=\sum_{D\geq 0}\tilde{R}_{L}(D)T^{\deg(D)}.$
The following result is crucial for the study of the series $\operatorname{Z}_{U,L}(s)$, as it allows us to relate this series to the zeta function of the base field $K$.
Lemma 5.14.
The following relation holds:
$\operatorname{Z}_{U,L}(T)=\left(\sum_{D\geq 0}R_{L}(D)T^{\deg(D)}\right)\frac{1}{\operatorname{Z}_{K}(T)}.$
Proof.
Since $(R_{L},\tilde{R}_{L})$ is a $\mu$-couple, using Lemma 5.6 we get
$\begin{split}\operatorname{Z}_{U,L}(T)&=\sum_{D\geq 0}\tilde{R}_{L}(D)T^{\deg(D)}=\sum_{D\geq 0}\sum_{0\leq D^{\prime}\leq D}\mu(D-D^{\prime})R_{L}(D^{\prime})T^{\deg(D)}\\ &=\sum_{D\geq 0}\sum_{D^{\prime}\geq 0}\mu(D)R_{L}(D^{\prime})T^{\deg(D+D^{\prime})}=\left(\sum_{D\geq 0}R_{L}(D)T^{\deg(D)}\right)\left(\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D^{\prime})}\right)\\ &=\left(\sum_{D\geq 0}R_{L}(D)T^{\deg(D)}\right)\frac{1}{\operatorname{Z}_{K}(T)},\end{split}$
where in the last equality we used Lemma 5.8. This proves the desired identity. ∎
5.3. Proof of Theorem 5.2
In this section we give the proof of Theorem 5.2, using the preliminary lemmas from Section 5.2.
For the sake of the reader’s convenience, we divide the proof into four independent steps.
Step 1. Rewriting the function $R_{L}(D)$. For positive integers $n,m$ and $D\in\operatorname{Div}^{+}(K)$, we define
$N_{m}^{n}(D):=\#\left\{(x_{i})\in\mathbb{A}^{m}:n\sup_{i\in I_{m}}(x_{i})_{\infty}\leq D\right\},$
and
$\tilde{N}_{m}(D):=\#\left\{(x_{i})\in\mathbb{A}^{m}:\sup_{i\in I_{m}}(x_{i})_{\infty}=D\right\}.$
Note that $(N_{m}^{1},\tilde{N}_{m})$ forms a $\mu$-couple and $N_{m}^{1}(D)=(N_{1}^{1}(D))^{m}=q^{m\ell(D)}$ (see Section 2.2).
For each integer $c\geq 1$, we also define the set
$\operatorname{Div}_{c}:=\{D\in\operatorname{Div}^{+}(K):v(D)\leq c\text{ for all }v\in\operatorname{Val}(K)\},$
and note that by the division algorithm all divisor $D\geq 0$ can be written uniquely as $D=(c+1)D_{1}+D_{2}$, where $D_{1}\geq 0$ and $D_{2}\in\operatorname{Div}_{c}$. Observe also that
$N_{1}^{c+1}((c+1)D_{1}+D_{2})=N_{1}^{1}(D_{1})\quad\text{for all }D_{1}\geq 0,D_{2}\in\operatorname{Div}_{c}.$ (17)
In particular, choosing $c=c_{L}-1$ where
$c_{L}:=\gamma a_{r}+\xi,$
we have that every $D\geq 0$ may be written uniquely as
$D=c_{L}D_{1}+D_{2},\quad\text{with }D_{1}\geq 0,D_{2}\in\operatorname{Div}_{c_{L}-1},$ (18)
and for every $D^{\prime}\geq 0$, we have
$c_{L}D^{\prime}\leq D\Leftrightarrow D^{\prime}\leq D_{1}.$ (19)
Now, recalling that $d=r+t-1$, it follows from the definition of $R_{L}(D)$ and $\mathbf{d}_{L}(x)$ that
$\begin{split}R_{L}(D)&=\#\left\{(x_{i})\in\mathbb{A}^{d}:\gamma\sup_{\begin{subarray}{c}i\in I_{t-1}\\ j\in I_{r-1}\end{subarray}}\big{(}(x_{d})_{\infty},(x_{t-1+j})_{\infty}+a_{j}(x_{i})_{\infty},a_{r}(x_{i})_{\infty}\big{)}+\xi\sup_{i\in I_{t-1}}\big{(}(x_{i})_{\infty}\big{)}\leq D\right\}\\ &=\sum_{0\leq D^{\prime}}\tilde{N}_{t-1}(D^{\prime})\#\big{\{}(x_{i})\in\mathbb{A}^{r}:\gamma\sup_{j\in I_{r-1}}((x_{r})_{\infty},(x_{j})_{\infty}+a_{j}D^{\prime},a_{r}D^{\prime})+\xi D^{\prime}\leq D\big{\}}.\end{split}$
Since
$c_{L}D^{\prime}\leq\gamma\sup_{j\in I_{r-1}}((x_{r})_{\infty},(x_{j})_{\infty}+a_{j}D^{\prime},a_{r}D^{\prime})+\xi D^{\prime},$
writing $D$ as in (18) we get (using (19))
$\begin{split}R_{L}(D)&=\sum_{0\leq D^{\prime}\leq D_{1}}\tilde{N}_{t-1}(D^{\prime})\#\big{\{}(x_{i})\in\mathbb{A}^{r}:\gamma\sup_{j\in I_{r-1}}((x_{r})_{\infty},(x_{j})_{\infty}+a_{j}D^{\prime},a_{r}D^{\prime})+\xi D^{\prime}\leq D\big{\}}.\end{split}$
Furthermore, using (19) we see that for $D^{\prime}$ satisfying $0\leq D^{\prime}\leq D_{1}$, the condition
$\gamma\sup_{j\in I_{r-1}}((x_{r})_{\infty},(x_{j})_{\infty}+a_{j}D^{\prime},a_{r}D^{\prime})+\xi D^{\prime}\leq D,$
is equivalent to the system of inequalities
$\left\{\begin{array}[]{ll}\gamma(x_{r})_{\infty}\leq D-\xi D^{\prime},&\\ \gamma(x_{j})_{\infty}\leq D-\left(\xi+\gamma a_{j}\right)D^{\prime},&\text{for all }j\in I_{r-1}.\end{array}\right.$
Therefore, we obtain
$R_{L}(D)=\sum_{0\leq D^{\prime}\leq D_{1}}\tilde{N}_{t-1}(D^{\prime})N_{1}^{\gamma}\left(D-\xi D^{\prime}\right)\prod_{j=1}^{r-1}N_{1}^{\gamma}\left(D-\left(\xi+\gamma a_{j}\right)D^{\prime}\right).$
Putting $a_{0}:=0$, we rewrite this identity as
$R_{L}(D)=\sum_{0\leq D^{\prime}\leq D_{1}}\tilde{N}_{t-1}(D^{\prime})\prod_{j=0}^{r-1}N_{1}^{\gamma}\left(D-\left(\xi+\gamma a_{j}\right)D^{\prime}\right).$ (20)
Step 2. Analysis of the formal series $\sum_{D\geq 0}R_{L}(D)T^{\deg(D)}$. Let us define
$\operatorname{Z}_{1}(T):=\sum_{D\geq 0}R_{L}(D)T^{\deg(D)}.$
Using (18) we have
$\operatorname{Z}_{1}(T)=\sum_{D_{1}\geq 0,D_{2}\in\operatorname{Div}_{c_{L}-1}}R_{L}\left(c_{L}D_{1}+D_{2}\right)T^{\deg\left(c_{L}D_{1}+D_{2}\right)}.$
Fixing now $D_{1}\geq 0,D_{2}\in\operatorname{Div}_{c_{L}-1}$ and using (20) we have
$R_{L}\left(c_{L}D_{1}+D_{2}\right)T^{\deg(c_{L}D_{1}+D_{2})}$
$=$ $\sum_{0\leq D^{\prime}\leq D_{1}}\tilde{N}_{t-1}(D^{\prime})\prod_{j=0}^{r-1}N_{1}^{\gamma}\left(c_{L}D_{1}+D_{2}-\left(\xi+\gamma a_{j}\right)D^{\prime}\right)T^{\deg\left(c_{L}D_{1}+D_{2}\right)}$
$=$ $\sum_{0\leq D^{\prime}\leq D_{1}}\tilde{N}_{t-1}(D^{\prime})\prod_{j=0}^{r-1}N_{1}^{\gamma}\left(c_{L}(D_{1}-D^{\prime})+D_{2}+\gamma(a_{r}-a_{j})D^{\prime}\right)T^{\deg\left(c_{L}(D_{1}-D^{\prime})+D_{2}+c_{L}D^{\prime}\right)}.$
Writing $D_{1}-D^{\prime}=D^{\prime\prime}$ with $D^{\prime\prime}\geq 0$ we get (using (17))
$\operatorname{Z}_{1}(T)$ $=\sum_{\begin{subarray}{c}D^{\prime}\geq 0,D^{\prime\prime}\geq 0\\ D_{2}\in\operatorname{Div}_{c_{L}-1}\end{subarray}}\tilde{N}_{t-1}(D^{\prime})\prod_{j=0}^{r-1}N_{1}^{\gamma}\left(c_{L}D^{\prime\prime}+D_{2}+\gamma(a_{r}-a_{j})D^{\prime}\right)T^{\deg\left(c_{L}D^{\prime\prime}+D_{2}+c_{L}D^{\prime}\right)}$
$=\sum_{D\geq 0,D^{\prime}\geq 0}\tilde{N}_{t-1}(D^{\prime})\prod_{j=0}^{r-1}N_{1}^{\gamma}\left(D+\gamma(a_{r}-a_{j})D^{\prime}\right)T^{\deg\left(D+c_{L}D^{\prime}\right)}$
$=\sum_{\begin{subarray}{c}D\geq 0,D^{\prime}\geq 0\\ D_{1}\in\operatorname{Div}_{\gamma-1}\end{subarray}}\tilde{N}_{t-1}(D^{\prime})\prod_{j=0}^{r-1}N_{1}^{\gamma}\left(\gamma D+D_{1}+\gamma(a_{r}-a_{j})D^{\prime}\right)T^{\deg\left(\gamma D+D_{1}+c_{L}D^{\prime}\right)}$
$=\sum_{\begin{subarray}{c}D\geq 0,D^{\prime}\geq 0\\ D_{1}\in\operatorname{Div}_{\gamma-1}\end{subarray}}\tilde{N}_{t-1}(D^{\prime})\prod_{j=0}^{r-1}N_{1}^{1}\left(D+(a_{r}-a_{j})D^{\prime}\right)T^{\deg\left(\gamma D+D_{1}+c_{L}D^{\prime}\right)}.$
Recall that $N_{X}=\#\{j\in\{1,\ldots,r\}:a_{j}=a_{r}\}$ and note that $j=r-N_{X}$ is the largest non-negative index satisfying $a_{j}<a_{r}$. Hence, we have
$\operatorname{Z}_{1}(T)$ $=\sum_{\begin{subarray}{c}D\geq 0,D^{\prime}\geq 0\\ D_{1}\in\operatorname{Div}_{\gamma-1}\end{subarray}}\tilde{N}_{t-1}(D^{\prime})N_{1}^{1}(D)^{N_{X}-1}\prod_{j=0}^{r-N_{X}}N_{1}^{1}\left(D+(a_{r}-a_{j})D^{\prime}\right)T^{\deg\left(\gamma D+D_{1}+c_{L}D^{\prime}\right)}.$
Now, define
$\operatorname{Z}_{2}(T)$ $:=\sum_{\begin{subarray}{c}D\geq 0,D^{\prime}\geq 0\\ D_{1}\in\operatorname{Div}_{\gamma-1}\end{subarray}}\tilde{N}_{t-1}(D^{\prime})N_{1}^{1}(D)^{N_{X}-1}\prod_{j=0}^{r-N_{X}}q^{1-g}q^{\deg(D)+(a_{r}-a_{j})\deg(D^{\prime})}T^{\deg\left(\gamma D+D_{1}+c_{L}D^{\prime}\right)}.$
Putting $A:=\min\{a_{r}-a_{j}:j\in\{0,\ldots,r-N_{X}\}\}=a_{r}-a_{r-N_{X}}$, we note by the Riemann–Roch theorem (see Section 2.1) that
$N_{1}^{1}\left(D+(a_{r}-a_{j})D^{\prime}\right)=q^{1-g}q^{\deg(D)+(a_{r}-a_{j})\deg(D^{\prime})}\quad\text{for all }j\in\{0,\ldots,r-N_{X}\},$
provided $\deg(D+AD^{\prime})>2g-2$. This implies
$\operatorname{Z}_{1}(T)-\operatorname{Z}_{2}(T)$ $=$ $\sum_{\begin{subarray}{c}D\geq 0,D^{\prime}\geq 0\\ \deg(D)+A\deg(D^{\prime})\leq 2g-2\\ D_{1}\in\operatorname{Div}_{\gamma-1}\end{subarray}}\tilde{N}_{t-1}(D^{\prime})N_{1}^{1}(D)^{N_{X}-1}\left(\prod_{j=0}^{r-N_{X}}N_{1}^{1}(D+(a_{r}-a_{j})D^{\prime})\right.$ (21)
$\hskip 56.9055pt\left.-q^{(r+1-N_{X})(1-g)}\prod_{j=0}^{r-N_{X}}q^{\deg(D)+(a_{r}-a_{j})\deg(D^{\prime})}\right)T^{\deg\left(\gamma D+D_{1}+c_{L}D^{\prime}\right)}$
$=$ $P_{1}(T)\sum_{D_{1}\in\operatorname{Div}_{\gamma-1}}T^{\deg(D_{1})},$
where $P_{1}(T)\in\mathbb{Q}[T]$, since $A>0$ and there are only finitely many pairs of divisors $D\geq 0,D^{\prime}\geq 0$ satisfying $\deg(D+AD^{\prime})\leq 2g-2$. Next, we further compute
$\operatorname{Z}_{2}(T)=$ $q^{(r+1-N_{X})(1-g)}\left(\sum_{D^{\prime}\geq 0}\tilde{N}_{t-1}(D^{\prime})q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg\left(D^{\prime}\right)}T^{c_{L}\deg(D^{\prime})}\right)$
×$\hskip 56.9055pt\times\left(\sum_{D\geq 0}N_{1}^{1}(D)^{N_{X}-1}\left(q^{r+1-N_{X}}T^{\gamma}\right)^{\deg(D)}\right)\left(\sum_{D_{1}\in\operatorname{Div}_{\gamma-1}}T^{\deg(D_{1})}\right).$
Putting
$\operatorname{Z}_{3}(T):=$ $q^{(r+1-N_{X})(1-g)}\left(\sum_{D^{\prime}\geq 0}\tilde{N}_{t-1}(D^{\prime})q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg\left(D^{\prime}\right)}T^{c_{L}\deg(D^{\prime})}\right)$
×$\hskip 56.9055pt\times\left(\sum_{D\geq 0}\left(q^{1-g+\deg(D)}\right)^{N_{X}-1}\left(q^{r+1-N_{X}}T^{\gamma}\right)^{\deg(D)}\right)\left(\sum_{D_{1}\in\operatorname{Div}_{\gamma-1}}T^{\deg(D_{1})}\right)$
$=$ $q^{r(1-g)}\left(\sum_{D^{\prime}\geq 0}\tilde{N}_{t-1}(D^{\prime})q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg\left(D^{\prime}\right)}T^{c_{L}\deg(D^{\prime})}\right)$
×$\hskip 56.9055pt\times\left(\sum_{D\geq 0}\left(q^{r}T^{\gamma}\right)^{\deg(D)}\right)\left(\sum_{D_{1}\in\operatorname{Div}_{\gamma-1}}T^{\deg(D_{1})}\right)$
we have, using the Riemann–Roch theorem once more, that
$\begin{split}\operatorname{Z}_{2}(T)-\operatorname{Z}_{3}(T)=&q^{(r+1-N_{X})(1-g)}\left(\sum_{D^{\prime}\geq 0}\tilde{N}_{t-1}(D^{\prime})q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg\left(D^{\prime}\right)}T^{c_{L}\deg(D^{\prime})}\right)\\ &\hskip 56.9055pt\times P_{2}(T)\left(\sum_{D_{1}\in\operatorname{Div}_{\gamma-1}}T^{\deg(D_{1})}\right),\end{split}$ (22)
with
$P_{2}(T):=\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}\left(N_{1}^{1}(D)^{N_{X}-1}-\big{(}q^{1-g+\deg(D)}\big{)}^{N_{X}-1}\right)\left(q^{r+1-N_{X}}T^{\gamma}\right)^{\deg(D)}\in\mathbb{Q}[T].$
Step 3. Rewriting the formal series $\operatorname{Z}_{3}(T)$. Taking into account that $(N_{t-1}^{1},\tilde{N}_{t-1})$ is a $\mu$-couple, and using Lemma 5.14, we have
$\begin{split}&\sum_{D^{\prime}\geq 0}\tilde{N}_{t-1}(D^{\prime})q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg\left(D^{\prime}\right)}T^{c_{L}\deg(D^{\prime})}\\ =&\sum_{D\geq 0}\sum_{0\leq D^{\prime}\leq D}\mu(D-D^{\prime})N_{t-1}^{1}(D^{\prime})q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg\left(D\right)}T^{c_{L}\deg(D)}\\ =&\sum_{D\geq 0}\sum_{D^{\prime}\geq 0}\mu(D)N_{t-1}^{1}(D^{\prime})q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg(D+D^{\prime})+}T^{c_{L}\deg(D+D^{\prime})}\\ =&\left(\sum_{D^{\prime}\geq 0}N_{1}^{1}(D^{\prime})^{t-1}q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg(D^{\prime})}T^{c_{L}\deg(D^{\prime})}\right)\left(\sum_{D\geq 0}\mu(D)q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg(D)}T^{c_{L}\deg(D)}\right)\\ =&\left(\sum_{D^{\prime}\geq 0}N_{1}^{1}(D^{\prime})^{t-1}q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg(D^{\prime})}T^{c_{L}\deg(D^{\prime})}\right)\frac{1}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)}.\end{split}$
Using again the Riemann–Roch theorem, we obtain
$\begin{split}&\sum_{D^{\prime}\geq 0}\tilde{N}_{t-1}(D^{\prime})q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg\left(D^{\prime}\right)}T^{c_{L}\deg(D^{\prime})}\\ =&\left(\sum_{D^{\prime}\geq 0}q^{(t-1)(1-g)}q^{(t-1)\deg(D^{\prime})}q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg(D^{\prime})}T^{c_{L}\deg(D^{\prime})}+P_{3}(T)\right)\frac{1}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)}\\ =&\left(q^{(t-1)(1-g)}\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|+t-1}T^{c_{L}}\right)+P_{3}(T)\right)\frac{1}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)},\end{split}$
where
$\begin{split}P_{3}(T):=&\sum_{\begin{subarray}{c}D^{\prime}\geq 0\\ \deg(D^{\prime})\leq 2g-2\end{subarray}}\left(N_{1}^{1}(D^{\prime})^{t-1}-q^{(t-1)(1-g+\deg(D^{\prime}))}\right)q^{\left((r+1)a_{r}-|\mathbf{a}|\right)\deg(D^{\prime})}T^{c_{L}\deg(D^{\prime})}\in\mathbb{Q}[T].\end{split}$
We conclude
$\begin{split}\operatorname{Z}_{3}(T)&=\left(q^{(t-1)(1-g)}\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|+t-1}T^{c_{L}}\right)+P_{3}(T)\right)\\ &\hskip 85.35826pt\times\frac{q^{r(1-g)}\operatorname{Z}_{K}\left(q^{r}T^{\gamma}\right)}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)}\left(\sum_{D_{1}\in\operatorname{Div}_{\gamma-1}}T^{\deg(D_{1})}\right).\end{split}$ (23)
Note that, by (22), we also get
$\begin{split}\operatorname{Z}_{2}(T)-\operatorname{Z}_{3}(T)=&q^{(r+1-N_{X})(1-g)}\left(q^{(t-1)(1-g)}\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|+t-1}T^{c_{L}}\right)+P_{3}(T)\right)\\ &\hskip 56.9055pt\times\frac{P_{2}(T)}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)}\left(\sum_{D_{1}\in\operatorname{Div}_{\gamma-1}}T^{\deg(D_{1})}\right).\end{split}$ (24)
Step 4. Analytic behaviour of $\zeta_{U,D}(s)$. Since all $D\geq 0$ can be written as $D=\gamma D_{1}+D_{2},$ with $D_{1}\geq 0$ and $D_{2}\in\operatorname{Div}_{\gamma-1}$, we have
$\operatorname{Z}_{K}(T)=\left(\sum_{D\geq 0}T^{\gamma\deg(D)}\right)\left(\sum_{D\in\operatorname{Div}_{\gamma-1}}T^{\deg(D)}\right).$
This implies
$\sum_{D\in\operatorname{Div}_{\gamma-1}}T^{\deg(D)}=\frac{\operatorname{Z}_{K}(T)}{\operatorname{Z}_{K}(T^{\gamma})},$
and by Lemma 5.14 together with (21) we get
$\operatorname{Z}_{U,L}(T)=\frac{\operatorname{Z}_{1}(T)}{\operatorname{Z}_{K}(T)}=\frac{\operatorname{Z}_{2}(T)}{\operatorname{Z}_{K}(T)}+\frac{P_{1}(T)}{\operatorname{Z}_{K}(T^{\gamma})}.$
Then, using (24) we have
$\begin{split}\operatorname{Z}_{U,L}(T)=&\frac{\operatorname{Z}_{3}(T)}{\operatorname{Z}_{K}(T)}+\frac{q^{(d+1-N_{X})(1-g)}P_{2}(T)\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|+t-1}T^{c_{L}}\right)}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)\operatorname{Z}_{K}(T^{\gamma})}\\ &\hskip 56.9055pt+\frac{q^{(r+1-N_{X})(1-g)}P_{3}(T)P_{2}(T)}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)\operatorname{Z}_{K}(T^{\gamma})}+\frac{P_{1}(T)}{\operatorname{Z}_{K}(T^{\gamma})}\end{split}$
Finally, using (23) we get
$\begin{split}\operatorname{Z}_{U,L}(T)=&\frac{q^{d(1-g)}\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|+t-1}T^{c_{L}}\right)\operatorname{Z}_{K}\left(q^{r}T^{\gamma}\right)}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)\operatorname{Z}_{K}(T^{\gamma})}+\frac{q^{r(1-g)}P_{3}(T)\operatorname{Z}_{K}\left(q^{r}T^{\gamma}\right)}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)\operatorname{Z}_{K}(T^{\gamma})}\\ &+\frac{q^{(d+1-N_{X})(1-g)}P_{2}(T)\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|+t-1}T^{c_{L}}\right)}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)\operatorname{Z}_{K}(T^{\gamma})}+\frac{q^{(r+1-N_{X})(1-g)}P_{3}(T)P_{2}(T)}{\operatorname{Z}_{K}\left(q^{(r+1)a_{r}-|\mathbf{a}|}T^{c_{L}}\right)\operatorname{Z}_{K}(T^{\gamma})}\\ &\hskip 14.22636pt+\frac{P_{1}(T)}{\operatorname{Z}_{K}(T^{\gamma})}.\end{split}$
By Lemma 5.12 we conclude
$\zeta_{U,L}(s)$ $=\frac{q^{d(1-g)}\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)s-\left((r+1)a_{r}-|\mathbf{a}|+t-1\right)\right)\zeta_{K}(\gamma s-r)}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)s-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(\gamma s)}$
$\hskip 14.22636pt+\frac{q^{r(1-g)}P_{3}(q^{-s})\zeta_{K}(\gamma s-r)}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)s-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(\gamma s)}$
$\hskip 28.45274pt+\frac{q^{(d+1-N_{X})(1-g)}P_{2}(q^{-s})\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)s-\left((r+1)a_{r}-|\mathbf{a}|+t-1\right)\right)}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)s-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(\gamma s)}$
$\hskip 42.67912pt+\frac{q^{(r+1-N_{X})(1-g)}P_{3}(q^{-s})P_{2}(q^{-s})}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)s-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(\gamma s)}+\frac{P_{1}(q^{-s})}{\zeta_{K}(\gamma s)}.$
It follows from the properties of the zeta function $\zeta_{K}(s)$ (see Section 2.2) that $\zeta_{U,L}(s)$ is a rational function in $q^{-s}$. Moreover, using Corollary 5.9 and (13) we see that $\zeta_{U,L}(s)$ converges absolutely for
$\Re(s)>\max\{A_{L},B_{L}\}=a(L),$
with $A_{L},B_{L}$ defined in (12). Finally, recalling that $r\geq 1,t\geq 2$, we get the following properties:
  1. (1)
    If $A_{L}=B_{L}$, then $\zeta_{U,L}(s)$ has a pole of order 2 at $s=a(L)$ with
    $\begin{split}\lim_{s\to a(L)}(s-a(L))^{2}\zeta_{U,L}(s)&=\frac{q^{d(1-g)}}{\zeta_{K}(t)\zeta_{K}(r+1)}\frac{\big{(}\textup{Res}_{s=1}\zeta_{K}(s)\big{)}^{2}}{(\gamma+a_{r}\xi)\gamma}\\ &=\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(\gamma+a_{r}\xi)\gamma(q-1)^{2}\log(q)^{2}}.\end{split}$
  2. (2)
    If $A_{L}<B_{L}$, then $\zeta_{U,L}(s)$ satisfies
    $\begin{split}\lim_{s\to a(L)}(s-a(L))\zeta_{U,L}(s)&=\left(q^{d(1-g)}\zeta_{K}(\gamma B_{L}-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-B_{L}})\right)\\ &\hskip 56.9055pt\times\frac{\textup{Res}_{s=1}\zeta_{K}(s)}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma B_{L})(\gamma a_{r}+\xi)}\\ &=\left(q^{d(1-g)}\zeta_{K}(\gamma B_{L}-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-B_{L}})\right)\\ &\hskip 56.9055pt\times\frac{h_{K}q^{1-g}}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma B_{L})(\gamma a_{r}+\xi)(q-1)\log(q)}.\end{split}$
    From the definition of $P_{2}(T)$ and (14) we have
    $\begin{split}&q^{d(1-g)}\zeta_{K}(\gamma B_{L}-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-B_{L}})\\ =&q^{(d+1-N_{X})(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{(N_{X}-1)\ell(D)+(r+1-N_{X}-\gamma B_{L})\deg(D)}+q^{d(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-(\gamma B_{L}-r)\deg(D)}\\ =&q^{(d+1-N_{X})(1-g)}\mathbf{R}_{K}(1-N_{X},\gamma B_{L}-r+N_{X}-1).\end{split}$
    Since this value is positive, we conclude that $\zeta_{U,L}(s)$ has a simple pole at $s=a(L)$ in this case.
  3. (3)
    If $A_{L}>B_{L}$, then $\zeta_{U,L}(s)$ satisfies
    $\begin{split}\lim_{s\to a(L)}(s-a(L))\zeta_{U,L}(s)&=\left(q^{d(1-g)}\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)A_{L}-\left((r+1)a_{r}-|\mathbf{a}|+t-1\right)\right)+q^{r(1-g)}P_{3}(q^{-A_{L}})\right)\\ &\hskip 56.9055pt\times\frac{\textup{Res}_{s=1}\zeta_{K}(s)}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)A_{L}-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(r+1)\gamma}\\ &=\left(q^{d(1-g)}\zeta_{K}\left(A_{L}\xi+|\mathbf{a}|-t+1\right)+q^{r(1-g)}P_{3}(q^{-A_{L}})\right)\\ &\hskip 56.9055pt\times\frac{h_{K}q^{1-g}}{\zeta_{K}\left(A_{L}\xi+|\mathbf{a}|\right)\zeta_{K}(r+1)\gamma(q-1)\log(q)}.\end{split}$
    Using the definition of $P_{3}(T)$ and (14), we get
    $\begin{split}&q^{d(1-g)}\zeta_{K}\left(A_{L}\xi+|\mathbf{a}|-t+1\right)+q^{r(1-g)}P_{3}(q^{-A_{L}})\\ =&q^{r(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{\ell(D)(t-1)-\deg(D)\left(A_{L}\xi+|\mathbf{a}|\right)}+q^{d(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-\left(A_{L}\xi+|\mathbf{a}|-t+1\right)\deg(D)}\\ =&q^{r(1-g)}\mathbf{R}_{K}\left(1-t,A_{L}\xi+|\mathbf{a}|\right).\end{split}$
    As in the previous case, we conclude that $\zeta_{U,L}(s)$ has a simple pole at $s=a(L)$.
Since the condition $A_{L}=B_{L}$ (resp. $A_{L}<B_{L}$, $A_{L}>B_{L}$) is equivalent to $\xi=\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma$ (resp. $\xi<\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma$, $\xi>\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma$), the desired result follows from the properties stated above. This completes the proof of the theorem.
5.4. Decomposition of Hirzebruch–Kleinschmidt varieties
In this section we explain in detail the decomposition (5) of Hirzebruch–Kleinschmidt varieties mentioned in the introduction.
First, note that for $1\leq r^{\prime}\leq r$ and $2\leq t^{\prime}\leq t$, there is a natural embedding
$X_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})\hookrightarrow X_{d}(a_{1},\ldots,a_{r})$ (25)
given in homogeneous coordinates by
$([x_{0}:(x_{ij})_{i\in I_{t^{\prime}},j\in I_{r^{\prime}}}],[y_{1}:\ldots:y_{t^{\prime}}])\mapsto([x_{0}:(\tilde{x}_{ij})_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t^{\prime}}:0:\ldots:0]),$
where
$\tilde{x}_{ij}:=\left\{\begin{array}[]{ll}x_{ij}&\textup{if }i\leq t^{\prime}\text{ and }j\leq r^{\prime},\\ 0&\textup{otherwise}.\end{array}\right.$
In other words, we enlarge the matrix $(x_{ij})$ of size $t^{\prime}\times r^{\prime}$ to a matrix of size $t\times r$ by adding zero columns on the right and zero rows on the bottom, and similarly we enlarge $[y_{1}:\ldots:y_{t^{\prime}}]$ by adding zeroes on the right.
We also consider the natural embeddings
$\mathbb{P}^{r^{\prime}}\hookrightarrow X_{d}(a_{1},\ldots,a_{r}),\quad[z_{0}:\ldots:z_{r^{\prime}}]\mapsto([z_{0}:(\tilde{z}_{ij})_{i\in I_{t},j\in I_{r}}],[1:0:\ldots:0]),$ (26)
for $1\leq r^{\prime}\leq r$, where
$\tilde{z}_{ij}:=\left\{\begin{array}[]{ll}z_{j}&\textup{if }i=1,1\leq j\leq r^{\prime}\\ 0&\textup{otherwise},\end{array}\right.$
and
$\mathbb{P}^{t-1}\hookrightarrow X_{d}(a_{1},\ldots,a_{r}),\quad[y_{1}:\ldots:y_{t}]\mapsto([1:(0)_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t}]).$ (27)
Note that the images of $\mathbb{P}^{t-1}$ and $\mathbb{P}^{r}$ inside $X_{d}(a_{1},\ldots,a_{r})$ under these embeddings intersect exactly at the point $P_{0}:=([1:(0)_{i\in I_{t},j\in I_{r}}],[1:0:\ldots:0])$.
For $1\leq r^{\prime}\leq r$ and $2\leq t^{\prime}\leq t$, let us identify $U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})$ with its image under (25). Similarly, let us identify $\mathbb{P}^{r^{\prime}}$ and $\mathbb{P}^{t-1}$ with their images under (26) and (27), which we denote by $\mathbb{P}^{r^{\prime}}_{1}$ and $\mathbb{P}^{t-1}_{2}$, respectively.
Lemma 5.15.
Let $X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety. If $r\geq 2$, then we have the disjoint union decomposition
$X_{d}(a_{1},\ldots,a_{r})=X_{d-1}(a_{1},\ldots,a_{r-1})\sqcup(\mathbb{P}^{r}_{1}\setminus\mathbb{P}^{r-1}_{1})\sqcup\bigg{(}\bigsqcup_{2\leq t^{\prime}\leq t}U_{t^{\prime}+r-1}(a_{1},\ldots,a_{r})\bigg{)}.$ (28)
If $r=1$, then we have the disjoint union decomposition
$X_{d}(a_{1})=\mathbb{P}^{t-1}_{2}\sqcup(\mathbb{P}^{1}_{1}\setminus\{P_{0}\})\sqcup\bigg{(}\bigsqcup_{2\leq t^{\prime}\leq t}U_{t^{\prime}}(a_{1})\bigg{)},$ (29)
where $P_{0}=([1:(0)_{i\in I_{t},j\in I_{r}}],[1:0:\ldots:0])$. Moreover, given $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big, we have the following properties:
  1. (1)
    For all $1\leq r^{\prime}\leq r$, the restriction of $H_{L}$ to $\mathbb{P}^{r^{\prime}}_{1}\simeq\mathbb{P}^{r^{\prime}}$ corresponds to $H_{\mathbb{P}^{r^{\prime}}}^{\gamma}$.
  2. (2)
    The restriction of $H_{L}$ to $\mathbb{P}^{t-1}_{2}\simeq\mathbb{P}^{t-1}$ corresponds to $H_{\mathbb{P}^{t-1}}^{\xi}$.
  3. (3)
    For $1\leq r^{\prime}\leq r$ and $2\leq t^{\prime}\leq t$, the restriction of $H_{L}$ to $U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})$ equals $H_{L^{\prime}}$ where $L^{\prime}:=\gamma h^{\prime}+\xi f^{\prime}$ with $\{h^{\prime},f^{\prime}\}$ the basis of $\operatorname{Pic}(X_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}}))$ given in Proposition 3.2.
Proof.
Assume $r\geq 2$ and let $x=([x_{0}:(x_{ij})_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t}])\in X_{d}(a_{1},\ldots,a_{r})$. If $x_{ir}=0$ for all $i\in I_{t}$, then $x\in X_{d-1}(a_{1},\ldots,a_{r-1})$. Assume this is not the case, and let $t^{\prime}:=\max\{i\in I_{t}:y_{i}\neq 0\}$. From equation (8) with $n=t^{\prime}$ and $m=i>t^{\prime}$ we see that
$x_{ij}=0\text{ for all }i\in I_{t}\text{ with }i>t^{\prime}\text{ and all }j\in I_{r}.$
Assume $t^{\prime}=1$. Then, $x_{ij}=0$ for all $i\in I_{t}$ with $i>1$ and all $j\in I_{r}$. Since $x\not\in X_{d-1}(a_{1},\ldots,a_{r-1})$, we have $x_{1r}\neq 0$, and this implies that $x\in\mathbb{P}^{r}_{1}\setminus\mathbb{P}^{r-1}_{1}$. Finally, assume $t^{\prime}>1$. Since $y_{t^{\prime}}\neq 0$, having $x_{t^{\prime}r}=0$ would imply $x_{ir}=0$ also for all $i\neq t^{\prime}$ (because of equation (8) with $n=t^{\prime}$ and $j=r$), but this contradicts the fact that $x\not\in X_{d-1}(a_{1},\ldots,a_{r-1})$. So, $x_{t^{\prime}r}\neq 0$ and we have in this case that $x\in U_{t^{\prime}+r-1}(a_{1},\ldots,a_{r})$. This proves that
$X_{d}(a_{1},\ldots,a_{r})=X_{d-1}(a_{1},\ldots,a_{r-1})\cup(\mathbb{P}^{r}_{1}\setminus\mathbb{P}^{r-1}_{1})\cup\bigg{(}\bigcup_{2\leq t^{\prime}\leq t}U_{t^{\prime}+r-1}(a_{1},\ldots,a_{r})\bigg{)}.$
Moreover, these unions are disjoint by construction. This proves (28). The proof of the decomposition (29) in the case $r=1$ is completely analogous. Finally, the statements about the various restrictions of the height function $H_{L}$ follow directly from (11). This proves the lemma. ∎
Remark 5.16.
In practice, we identify the components $\mathbb{P}^{r}_{1}\setminus\mathbb{P}^{r-1}_{1}$ and $\mathbb{P}^{1}_{1}\setminus\{P_{0}\}$ in (28) and (29) with $\mathbb{A}^{r}$ and $\mathbb{A}^{1}$, respectively, as done in Example 1.4 in the introduction.
5.5. The anticanonical height zeta function - Proof of Theorem 1.2
We now prove Theorem 1.2 from the introduction. We choose $L=-K_{X}=\gamma f+\xi f$ with $\gamma=r+1$ and $\xi=t-|\mathbf{a}|$. Hence, $A_{L}=B_{L}=1$, $a(L)=1$ and $b(L)=2$. Assume $a_{r}>0$. Then, the absolute convergence of $\zeta_{U,-K_{X}}(s)$ for $\Re(s)>1$ and the fact that $\zeta_{U,-K_{X}}(s)$ is a rational function in $q^{-s}$ follow directly from Theorem 5.2. Moreover, we have
$\lim_{s\to 1}(s-1)^{2}\zeta_{U,-K_{X}}(s)=C$
with $C$ given by (4). This proves the result in the case $a_{r}>0$. Now, assume $a_{r}=0$. On the one hand, Theorem 5.4 shows that $\zeta_{X,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$, it is a rational function in $q^{-s}$ and
$\lim_{s\to 1}(s-1)^{2}\zeta_{X,-K_{X}}(s)=C$
with the same $C$ as before. On the other hand, by Lemma 5.15 we have
$X\setminus U\simeq\left\{\begin{array}[]{ll}X_{d-1}(0,\ldots,0)\sqcup(\mathbb{P}^{r}_{1}\setminus\mathbb{P}^{r-1}_{1})\sqcup\bigg{(}\bigsqcup_{2\leq t^{\prime}<t}U_{t^{\prime}+r-1}(a_{1},\ldots,a_{r})\bigg{)}&\text{if }r>1,\\ \mathbb{P}^{t-1}_{2}\sqcup(\mathbb{P}^{1}_{1}\setminus\{P_{0}\})\sqcup\bigg{(}\bigsqcup_{2\leq t^{\prime}<t}U_{t^{\prime}+r-1}(a_{1})\bigg{)}&\text{if }r=1,\end{array}\right.$
with $X_{d-1}(0,\ldots,0)\simeq\mathbb{P}^{t-1}\times\mathbb{P}^{r-1}$ in the case $r>1$. Using Theorems 5.4 and 4.1, we see that
$\zeta_{X\setminus U,-K_{X}}(s)=\sum_{P\in X(K)\setminus U(K)}H(P)^{-s}$
converges absolutely for $\Re(s)>1$, is a rational function in $q^{-s}$ and has a simple pole at $s=1$. We conclude that
$\zeta_{U,-K_{X}}(s)=\zeta_{X,-K_{X}}(s)-\zeta_{X\setminus U,-K_{X}}(s)$
satisfies the desired properties. This completes the proof of Theorem 1.2.
6. Example: Hirzebruch surfaces
In this last section we apply Theorems 4.1, 5.2 and 5.4 to the case of Hirzebruch surfaces $X:=X_{2}(a)$ with $a>0$ an integer, and general $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big. In this context, Lemma 5.15 gives
$X\simeq\mathbb{P}^{1}\sqcup\mathbb{A}^{1}\sqcup U_{2}(a),$ (30)
with associated height zeta functions $\zeta_{\mathbb{P}^{1}}(\xi s),\zeta_{\mathbb{A}^{1}}(\gamma s)=\zeta_{\mathbb{P}^{1}}(\gamma s)-1$ and $\zeta_{U,L}(s)$ where $U:=U_{2}(a)$. Moreover, we have
$A_{L}=\frac{2}{\gamma},\quad B_{L}=\frac{a+2}{\gamma a+\xi},$
and the analytic properties of $\zeta_{U,L}(s)$ can be extracted from Theorem 5.2. In particular, $\zeta_{U,L}(s)$ converges absolutely for $\Re(s)>\max\{A_{L},B_{L}\}$, it has a pole at $s=\max\{A_{L},B_{L}\}$, and this pole is of order 2 if $A_{L}=B_{L}$ and of order 1 otherwise. For the zeta function $\zeta_{\mathbb{A}^{1}}(\gamma s)$, we know by Theorem 4.1 that it converges absolutely for $\Re(s)>A_{L}$ and it has a simple pole at $s=A_{L}$. Finally, the zeta function $\zeta_{\mathbb{P}^{1}}(\xi s)$ has no finite abscissa of absolute convergence if $\xi\leq 0$, and it converges absolutely for $\Re(s)>\frac{2}{\xi}$ with a simple pole at $s=\frac{2}{\xi}$ if $\xi>0$. This allows for a complete analysis of the analytic properties of the height zeta function of each component in (30).
In the case $L=-K_{X}$, we have $\gamma=2,\xi=2-a$ and $A_{L}=B_{L}=1$. As in Example 1.4, we find it necessary to remove the component $\mathbb{P}^{1}$ in (30) in order to verify the analogue of Conjecture 1.1 over global function fields. Indeed, the zeta series $\zeta_{X\setminus\mathbb{P}^{1}}(s)=\zeta_{\mathbb{A}^{1}}(\gamma s)+\zeta_{U,L}(s)$ converges absolutely for $\Re(s)>1$, it is a rational function in $q^{-s}$ and satisfies
$\lim_{s\to 1}(s-1)^{2}\zeta_{X\setminus\mathbb{P}^{1}}(s)=\lim_{s\to 1}(s-1)^{2}\zeta_{U,L}(s)=\frac{q^{4(1-g)}h_{K}^{2}}{\zeta_{K}(2)^{2}(a+2)2(q-1)^{2}\log(q)^{2}}.$
This is in accordance with [Bou02, Proposition 3.1] (see also p. 355 in loc. cit.) where the case of the anticanonical height zeta function over the good open subset $U_{2}(a)$ was studied.
Let us now choose $X=X_{2}(1)$ (i.e. $a=1$) for simplicity in order to illustrate further our results for more general big classes $L=\gamma h+\xi f\in\operatorname{Pic}(X)$. In the following table we denote by $\sigma_{1},\sigma_{2}$ and $\sigma_{3}$ the abscissas of absolute convergence of $\zeta_{\mathbb{P}^{1}}(\xi s),\zeta_{\mathbb{A}^{1}}(\gamma s)$ and $\zeta_{U,L}(s)$, respectively, for the first possible choices of $\gamma>0$ and $\xi>-\gamma$. We also record the order $b(L)$ of the pole of $\zeta_{U,L}(s)$ at $s=\sigma_{3}$.
$\gamma$ $\xi$ $\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $b(L)$
1 0 $\infty$ 2 3 1
1 1 2 2 2 1
1 2 1 2 2 1
2 $-1$infty 1 3 1
2 0 $\infty$ 1 $3/2$ 1
2 1 2 1 1 2
3 $-2$infty $2/3$ 3 1
3 $-1$infty $2/3$ $3/2$ 1
3 0 $\infty$ $2/3$ $1$ 1
The case $\gamma=\xi=1$ seems particularly interesting, since $\sigma_{1}=\sigma_{2}=\sigma_{3}=2$ and $b(L)=1$, meaning that the three components in (30) have height zeta functions with a simple pole at $s=2$. Moreover, we can compute
$\lim_{s\to 2}(s-2)\zeta_{\mathbb{P}^{1}}(s)=\lim_{s\to 2}(s-2)\zeta_{\mathbb{A}^{1}}(s)=\frac{h_{K}q^{2(1-g)}}{\zeta_{K}(2)(q-1)\log(q)},$
and using that $\xi>\frac{1}{2}\gamma$ we get
$\lim_{s\to 2}(s-2)\zeta_{U,L}(s)=\frac{h_{K}q^{2(1-g)}\mathbf{R}_{K}\left(-1,3\right)}{\zeta_{K}(3)\zeta_{K}(2)(q-1)\log(q)}.$
In the special case $K=\mathbb{F}_{q}(T)$ we can use that $g=0,h_{K}=1$ and
$\zeta_{K}(s)=\frac{1}{(1-q^{1-s})(1-q^{-s})}$
(see e.g. [Ros02, p. 52]) and write
$\lim_{s\to 2}(s-2)\zeta_{\mathbb{P}^{1}}(s)=\lim_{s\to 2}(s-2)\zeta_{\mathbb{A}^{1}}(s)=\frac{q^{2}-1}{q\log(q)},$
and
$\lim_{s\to 2}(s-2)\zeta_{U,L}(s)=\frac{(q^{2}-1)(q^{2}+q+1)}{q^{2}\log(q)}.$
(see Remark 5.3(1)). Hence, in the case of the Hirzebruch surface $X_{2}(1)$ over $K=\mathbb{F}_{q}(T)$, with respect to the height function $H_{L}$ associated with $L=h+f\in\operatorname{Pic}(X_{2}(1))$, we can say that the component $U=U_{2}(1)$ in (30) has asymptotically more rational points of bounded height than the union $\mathbb{P}^{1}\sqcup\mathbb{A}^{1}$ of the other two components, since the residue of $\zeta_{U,L}(s)$ at $s=2$ is larger than the sum of the residues of $\zeta_{\mathbb{P}^{1}}(s)$ and $\zeta_{\mathbb{A}^{1}}(s)$.